### v2

# Linear Contextual Bandits with Hybrid Payoff: Revisited

Abstract
We study the Linear Contextual Bandit (LinearCB) problem in the hybrid reward setting. In this setting, every arm’s reward model contains arm specific parameters in addition to parameters shared across the reward models of all the arms. We can easily reduce this setting to two closely related settings; (a) Shared - no arm specific parameters, and (b) Disjoint - only arm specific parameters, enabling the application of two popular state of the art algorithms - LinUCB and DisLinUCB (proposed as Algorithm $1$ in Li et al. 2010). When the arm features are stochastic and satisfy a popular diversity condition, we provide new regret analyses for both LinUCB and DisLinUCB that significantly improves upon the known regret guarantees of these algorithms. Our novel analysis critically exploits the structure of the hybrid rewards and diversity of the arm features. Along with proving these new guarantees, we introduce a new algorithm HyLinUCB that crucially modifies LinUCB (using a new exploration coefficient) to account for sparsity in the hybrid setting. Under the same diversity assumptions, we prove that at the end of $T$ rounds, HyLinUCB also incurs only $\tilde{O}(\sqrt{T})$ regret. We perform extensive experiments on synthetic and real-world datasets demonstrating strong empirical performance of HyLinUCB. When the number of arm specific parameters is much larger than the number of shared parameters, we observe that DisLinUCB incurs the lowest regret. In this case, regret of HyLinUCB is the second best and it is extremely competitive to DisLinUCB. In all other situations, including our real-world dataset, HyLinUCB has significantly lower regret than LinUCB, DisLinUCB and other state of the art baselines we considered. We also empirically observe that the regret of HyLinUCB grows much slower with the number of arms $K$, compared to baselines, making it suitable even for very large action spaces.
Keywords Linear Contextual Bandits, Hybrid Payoff, Stochastic Contexts
1 Introduction
The LinearCB problem is a popular choice to model sequential decision making scenarios such as news recommendations (Li et al., 2010), web-page optimization (Hill et al., 2017) etc. In LinearCB, at each round, a learner receives (from the environment) a set of arms (actions) specified as feature vectors, selects one of them, and receives a reward sampled from a linear model (over the features of the selected arm). The goal of the learner is to design a policy of selecting arms that minimizes its cumulative regret. Here, regret of a round is calculated as the difference between the best possible reward in that round and the actual reward received.
The LinearCB problem is most commonly studied under the shared setting. Here, the linear reward model is shared across all the arms and algorithms such as LinUCB (Li et al., 2010; Lattimore and Szepesvári, 2020) achieve state of the art regret. However, this setting can be quite restrictive in applications where the reward model is different for different arms. For example, in news recommendation, where arms are different news items, the reward model (say click through rate) for news items about sports could be very different from that of politics, movies etc. As a result, the problem has also been studied under the disjoint and hybrid settings Li et al. (2010). While in the disjoint setting, the reward model of each arm only contains parameters specific to the arm (called arm-specific parameters), in the hybrid setting they also contain parameters shared across all the arms (called shared parameters). Algorithms were developed for the disjoint and hybrid settings in Algorithm $1$ and Algorithm $2$ (respectively) of Li et al. (2010). While Algorithm $1$ (called DisLinUCB from here on-wards) can be analyzed using ideas similar to the analysis of LinUCB for the shared setting (Lattimore and Szepesvári, 2020), Algorithm $2$ requires tuning of a hyper-parameter. Analyzing the regret of this algorithm is quite non-trivial and to the best of our knowledge tight regret guarantees for the hybrid setting are not known. Designing algorithms for the hybrid setting that overcome this challenge and also have strong regret guarantees is the main focus of our work. For each arm’s reward model, let $d_{1},d_{2}$ be the number of shared and arm specific parameters respectively, and let $K$ be the total number of arms. Further, define $d\coloneqq d_{1}+d_{2}K$. We make the following contributions.
1.1 Our Contributions
  1. 1.
    First, we reduce the hybrid setting to the shared setting with $d$ parameters and provide a new analysis for LinUCB (Lattimore and Szepesvári, 2020). We prove that if features of arms pulled by LinUCB satisfy Assumption 1, then, at the end of $T$ rounds, LinUCB incurs a regret of $\tilde{O}(\sqrt{dKT})$, when $T=\tilde{\Omega}(K^{3})$. Note that the standard guarantee of LinUCB for this reduced problem is $\tilde{O}(d\sqrt{T})$.
  2. 2.
    Next, by reducing to the disjoint setting, we provide a new analysis for DisLinUCB. We prove that, under the same assumption as above, at the end of $T$ rounds, DisLinUCB incurs a regret of $\tilde{O}\left(\sqrt{(d_{1}+d_{2})KT}\right)$, when $T=\tilde{\Omega}(1)$. The standard analysis (following analysis of LinUCB in Lattimore and Szepesvári (2020)) implies $\tilde{O}((d_{1}+d_{2})\sqrt{KT})$ regret, which is much worse.
  3. 3.
    Finally, we modify LinUCB using a tighter exploration parameter and develop a new algorithm HyLinUCB. Under the same diversity assumption, we prove a $\tilde{O}(\sqrt{K^{3}T}+\sqrt{dKT})$ regret guarantee for this algorithm. While our guarantee has a weaker dependence on $K$ compared to the above algorithms, we empirically observe it to be much stronger compared to them. By performing extensive experiments on synthetic (capturing a wide range of problem settings) and real-world datasets (Yahoo! Front Page Dataset Yahoo! Webscope (2010)), we demonstrate that in almost all cases HyLinUCB has much lower regret than LinUCB, DisLinUCB and other baselines. When the number of shared parameters is larger than the arm-specific parameters, regret of HyLinUCB is significantly lower than all state of the art baselines we compare it with. In the other case, i.e., when the number of arm-specific parameters are larger, HyLinUCB is extremely competitive with the best algorithm i.e. DisLinUCB. We also empirically study the variation in regret with respect to the number of arms $K$, and observe that regret of HyLinUCB grows much slower with $K$, compared to baselines, making it suitable even for very large action spaces.
1.2 Additional Remarks on Contributions
We make some additional remarks to provide more clarity on our contributions.
Reduction to the shared setting: The hybrid setting is easily reduced to the shared setting by combining the parameters of linear reward models of all the actions into a common reward model. We provide complete details of this reduction in Section 2. This reduction helps us apply most of the known algorithms for LinearCB (discussed in detail in Section 1.3). Note that the reduction significantly increases the dimensionality of the arm features and the known regret guarantees of these LinearCB algorithms might not be optimal anymore. The additional structure in the problem needs to be exploited in the analysis to improve these guarantees. Our first contribution does this for LinUCB. Similar improved analysis might exist for other state of the art LinearCB algorithms we discuss in Section 1.3. We leave this problem for future work.
Regret guarantee of HyLinUCB: Even though our regret guarantee for HyLinUCB has a slightly worse dependence on $K$, we believe this is an artifact of our proof and that it can be improved further. Our synthetic and real-data experiments provide strong evidence of this as HyLinUCB performs much better than LinUCB in all the different problem settings we consider. Regret guarantee that improves this dependence is an interesting direction for future work.
Diversity Assumption: Our regret guarantees are derived under the assumption that the arms pulled by the algorithms satisfy a diversity condition (Assumption 1) that has been studied in (Chatterji et al., 2020; Ghosh et al., 2021; Gentile et al., 2017). The assumption states that the minimum eigenvalue of the expected outer product of the pulled arm’s feature vector with itself, is bounded away from zero. This assumption allows the algorithms to perform parameter estimation in conjunction to regret minimization, the former being crucial to derive better dependence on the number of arms in the regret upper bound. The necessity of such diversity assumptions and their implications on better regret rates has been studied in (Papini et al., 2021). Even though the assumption is algorithm dependent, there are problem instances for which all algorithms satisfy it (Section 3, (Papini et al., 2021)). We empirically show that Assumption 1 is indeed satisfied by the algorithms we study in this work.
1.3 Related Work
As highlighted in Section 1.2, the LinearCB problem in the hybrid setting can be reduced to the shared setting, enabling the application of all LinearCB algorithms designed for the shared setting. Over the last couple of decades, there has been a substantial progress on LinearCB for the shared setting. Many state of the art algorithms follow the optimism in the face of uncertainty approach (Lai and Robbins, 1985), and provide regret analysis with near optimal guarantees (Auer, 2003; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2008; Chu et al., 2011; Abbasi-Yadkori et al., 2011). For general stochastic arm features, the best known regret upper bound (independent of number of arms $K$) provided in these works is $\tilde{O}(d\sqrt{T})$ and the best known lower bound (Chu et al., 2011) is $\tilde{\Omega}(d\sqrt{T})$, where $d$ is the dimensionality of the arm features and $T$ is the number of rounds. The LinUCB algorithm (Li et al., 2010) attains this upper bound ($\tilde{O}(d\sqrt{T})$) Lattimore and Szepesvári (2020) and is also a popular choice for applications such as news recommendation (Li et al., 2010) due to it’s strong empirical performance. Moreover, it was also adapted to the hybrid setting in Li et al. (2010). As a result we choose it as one of our main candidate algorithms to study in this paper. The popular SupLinUCB (Chu et al., 2011) algorithm improves the dependence on $d$ and guarantees $\tilde{O}(\sqrt{dT}\log^{3/2}{K})$ regret where $K$ is the number of actions. A recent variant of the SupLinUCB algorithm (Li et al., 2019) improves this further to $\tilde{O}(\sqrt{dT}\log{K})$. While the dependence on $d$ improves, it is traded off with a logarithmic dependence on $K$. It has also been noted (Kim et al., 2023) that despite strong regret guarantees, SupLinUCB and its variants (Li et al., 2017; 2019) explore excessively and perform worse in practice due to computational inefficiency. Recently, Kim et al. (2023) developed a new algorithm HyRan which does not depend on $K$ in the leading terms and attains a regret guarantee of $\tilde{O}(\sqrt{dT})$ (improving significantly over SupLinUCB both theoretically and empirically). While they prove this under an additional assumption on stochasticity of the arm features, they also provide a matching lower bound showing tightness of their guarantee. Their experiments validate their strong upper bound by demonstrating superior performance compared to many state of the art LinearCB baselines for a large variety of problem instances. Due to their strong guarantee and empirical performance, we compare our algorithms with HyRan in Section 4. Another family of algorithms (based on randomized exploration) referred to as Thompson Sampling (Thompson, 1933) based methods are an active area of research. Linear contextual version of this method was recently developed (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017) and best known regret guarantees are either $\tilde{O}(d\sqrt{T\log{K}})$ or $\tilde{O}(d^{3/2}\sqrt{T})$. Both guarantees are worse than that of LinUCB. Even though Thompson Sampling does well practically, the HyRan algorithm from Kim et al. (2023) was shown to be superior to them for all problem instances considered in Kim et al. (2023). Since we already compare with HyRan in our experiments in Section 4, we do not compare to Thompson Sampling based methods.
2 Problem Formulation
[…]
LinearCB with hybrid rewards: We assume there exist unknown parameter vectors $\theta^{*}\in\mathbb{R}^{d_{1}}$ and $\beta_{i}^{*}\in\mathbb{R}^{d_{2}}$ for $i\in[K]$. At each round $t\in\mathbb{N}$, the learner receives a set of $K$ feature vector tuples $\mathcal{X}_{t}=\{(x_{i,t},z_{i,t}):i\in[K]\}$. For each arm $i\in[K]$, the features $x_{i,t}\in\mathbb{R}^{d_{1}}$ correspond to the parameter vector $\theta^{*}$ and the features $z_{i,t}\in\mathbb{R}^{d_{2}}$ correspond to the arm-specific parameters $\beta_{i}^{*}$, i.e. if the learner selects arm $i_{t}\in[K]$ at round $t$, she receives a reward
$r_{t}=\langle x_{i_{t},t},\theta^{*}\rangle+\langle z_{i_{t},t},\beta_{i_{t}}^{*}\rangle+\eta_{t},$
where $\eta_{t}$ is a conditionally $1$-subgaussian random noise. Specifically, let $\mathcal{F}_{t}$ be the filtration $(\mathcal{X}_{1},i_{1},r_{1},\dots,r_{t-1})$, then, $\mathbb{E}[\eta_{t}\mid\mathcal{F}_{t}]=0,\text{ and, }\mathbb{E}[e^{\alpha\eta_{t}}\mid\mathcal{F}_{t}]\leq e^{\alpha^{2}/2}$, for all $\alpha\in\mathbb{R}$. Note that we do not assume the noise variable $\eta_{t}$ to depend on the arms selected by the learner. This later allows us to reduce the problem to the shared setting albeit in a larger dimension. We define the (cumulative) regret of the learner Alg at the end of round $T$ as,
$\texttt{Reg}(T,\texttt{Alg})=$ $\sum\limits_{t=1}^{T}\max_{j\in[K]}\left(\langle{x_{j,t}},{\theta^{*}}\rangle+\langle{z_{j,t}},{\beta_{j}^{*}}\rangle\right)-\sum\limits_{t=1}^{T}\left(\langle{x_{i_{t},t}},{\theta^{*}}\rangle+\langle{z_{i_{t},t}},{\beta_{i_{t}}^{*}}\rangle\right).$
Assumptions: The key assumption in this work is the diversity assumption, similar to the one made in (Chatterji et al., 2020; Ghosh et al., 2021; Gentile et al., 2017) for the shared setting. We extend this assumption (stated below) to the hybrid setting by adding one extra assumption that arises from the hybrid nature of the problem.
Assumption 1 (Diversity).
There exists a constant $\rho>0$, such that features of the arm selected by the algorithm, i.e., $x_{i_{t},t}$, $z_{i_{t},t}$, for all rounds $t\in\mathbb{N}$, satisfy,
  1. 1.
    $\mathbb{E}[x_{i_{t},t}\mid\mathcal{F}_{t-1}]={\mathbf{0}},\mathbb{E}[z_{i_{t},t}\mid\mathcal{F}_{t-1}]={\mathbf{0}},$ and $\mathbb{E}\left[x_{i_{t},t}z_{i_{t},t}^{\intercal}\mid\mathcal{F}_{t-1}\right]={\mathbf{0}}$,
  2. 2.
    $\mathbb{E}\left[x_{i_{t},t}x_{i_{t},t}^{\intercal}\mid\mathcal{F}_{t-1}\right]\succcurlyeq\rho\mathbf{I}_{d_{1}},\mathbb{E}\left[z_{i_{t},t}z_{i_{t},t}^{\intercal}\mid\mathcal{F}_{t-1}\right]\succcurlyeq\rho\mathbf{I}_{d_{2}}$
We note that this assumption is algorithm dependent, however, there are problem instances in the shared setting for which any algorithm satisfies it (Section $3$, (Papini et al., 2021)). We empirically validate this assumption in the hybrid setting for the algorithms studied in this work, thereby demonstrating that the assumption and its crucial implications indeed hold in practice. The details of this empirical study is presented in Appendix F. We also make the following standard assumption on the reward parameters.
Assumption 2.
There is a fixed constant $S\in\mathbb{R}$ (known to us) such that the reward parameters $\left\lVert\theta^{*}\right\rVert_{2}\leq S$ and $\left\lVert\beta^{*}_{i}\right\rVert_{2}\leq S$ for all $i\in[K]$. Further, $\left\lVert x_{i,t}\right\rVert_{2}\leq 1$ and $\left\lVert z_{i,t}\right\rVert_{2}\leq 1$, for all arms $i\in[K]$ and rounds $t\in\mathbb{N}$.
Reduction to shared setting: We will now formalize our discussion from Section 1.2 about reducing the hybrid setting to the shared setting. We first embed the arm features into a sparse feature vector in $\mathbb{R}^{d_{1}+d_{2}K}$ via transformation $\mathcal{P}:[K]\times\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}^{d_{1}+d_{2}K}$ which takes as input arm $i\in[K]$ and its features $x\in\mathbb{R}^{d_{1}},z\in\mathbb{R}^{d_{2}}$ and maps it as follows:
$\mathcal{P}(i,x,z)=(x^{\intercal},0,\ldots,0,z^{\intercal},0,\ldots,0)^{\intercal}$ (1)
that is, the coordinate entries of $x$ are copied to the first $d_{1}$ coordinates of $\mathcal{P}(i,x,z)$ while the coordinate entries of $z$ are copied to the location starting from $d_{1}+(i-1)d_{2}+1$ and ending at $d_{1}+id_{2}$. We also combine the true shared and arm-specific parameters into a global parameter vector $\phi^{*}=(\theta^{*\intercal},\beta_{1}^{*\intercal},\ldots,\beta_{K}^{*\intercal})^{\intercal}$. It’s easy to see that this transformation converts the hybrid setting to the shared setting while preserving the mean reward of every arm, i.e.,
$\langle{\mathcal{P}(i,x,z)},{\phi^{*}}\rangle=\langle{x},{\theta^{*}}\rangle+\langle{z},{\beta^{*}_{i}}\rangle\quad\forall\ i\in[K].$
Moreover, since the noise variable $\eta_{t}$ at round $t$ is assumed to be independent of the arm selected, the noisy rewards $r_{t}=\langle{\mathcal{P}(i,x,z)},{\phi^{*}}\rangle+\eta_{t}$ in this shared setting exactly captures the reward received by arm $i$ in the hybrid setting.
More Notations. For a matrix $\mathbf{M}\in\mathbb{R}^{p\times p}$ and vector $a\in\mathbb{R}^{p}$, for any $p>0$, we define $\left\lVert a\right\rVert_{\mathbf{M}}=\sqrt{a^{\intercal}\mathbf{M}a}$. For a square matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$, we will denote by $\lambda_{max}(\mathbf{A})$ and $\lambda_{min}(\mathbf{A})$ its maximum and minimum eigenvalues respectively. Finally, for $p,q\in\mathbb{R}$, we will write $p\vee q=\max\left\{p,q\right\}$ and $p\wedge q=\min\left\{p,q\right\}$.
3 Algorithms and Analysis
In this section, we present our main algorithms and their analyses. First, in Section 3.1 we present the LinUCB and DisLinUCB algorithms for the hybrid setting in Algorithms 1 and 2 respectively. These algorithms are well known, however, for completeness we present them with the minor modifications we need to make to apply them to the hybrid setting. Following this, in Theorem 3 and Corollary 4, we present the improved regret guarantees for LinUCB and DisLinUCB (under Assumptions 1 and 2) respectively. Next, in Section 3.2, we introduce the HyLinUCB algorithm which is exactly the same as LinUCB presented in Algorithm 1 but with a new confidence parameter $\gamma$ that we describe. We provide some intuition regarding our choice of this parameter and then in Theorem 5 present a regret guarantee for this algorithm under Assumptions 1 and 2.
3.1 LinUCB and DisLinUCB
[…]
Algorithm 1 LinUCB ($\lambda,\gamma$)
0:  Regularizer $\lambda$, exploration coefficient $\gamma$, failure probability $\delta\in(0,1)$
1:  Initialize $\widehat{\phi}={\mathbf{0}}\in\mathbb{R}^{d_{1}+d_{2}K}$, $\mathbf{M}=\lambda\mathbf{I}_{d_{1}+d_{2}K}$, $u={\mathbf{0}}\in\mathbb{R}^{d_{1}+d_{2}K}$
2:  for $t=1,\dots,T$ do
3:     Receive context $\mathcal{X}_{t}=\{(x_{i},z_{i})\}_{i\in[K]}$
4:     $\tilde{x}_{i}\coloneqq\mathcal{P}(i,x_{i},z_{i})$ for all $i\in[K]$
5:     $i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\tilde{x}_{i}},{\widehat{\phi}}\rangle+\gamma\left\lVert\tilde{x}_{i}\right\rVert_{\mathbf{M}^{-1}}$
6:     Play arm $i_{t}$ and observe reward $r_{t}$
7:     $u\leftarrow u+r_{t}\tilde{x}_{t}$
8:     $\mathbf{M}\leftarrow\mathbf{M}+\tilde{x}_{t}\tilde{x}_{t}^{\intercal}$
9:     Update: $\widehat{\phi}=\mathbf{M}^{-1}u$
10:  end for
Theorem 3 (Regret of LinUCB).
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)KT\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
The reduction from the hybrid setting to the shared setting leads to a sparse design matrix. Our proof of Theorem 3 critically exploits the structure of this sparse matrix and its inverse, along with some key technical ideas involving geometry of random vectors and eigenvalue bounds on random matrices applied to the stochastic arm features and the design matrix. For brevity, we present the complete proof in Appendix A.
Our second algorithm DisLinUCB presented in Algorithm 2 (same as Algorithm $1$ in (Li et al., 2010)) takes the same set of inputs as Algorithm 1 above, i.e., the total number of rounds $T$, a regularization parameter $\lambda$, an exploration coefficient $\gamma$ and a desired failure probability $\delta\in(0,1)$. In Step 1, for each arm $i\in[K]$, it performs initialization of the reward parameter vectors $\phi_{i}$ and $(d_{1}+d_{2})\times(d_{1}+d_{2})$ size design matrices $\mathbf{V}_{i}$. This algorithm treats the hybrid setting as a disjoint setting and therefore the dimensionality of the problem remains the same as the hybrid setting i.e. $d_{1}+d_{2}$. The algorithm runs for rounds $t=1$ to $t=T$ (Steps 2-9). In Steps 3,4, it receives the set $\mathcal{X}_{t}$ of arm features. For each arm $i\in[K]$, it concatenates the shared and disjoint arm features to obtain the $d_{1}+d_{2}$ dimensional feature vector $\overline{x}_{i}$ and treats it as the vector of arm-specific parameters in the algorithm. Then, in Step 5, it selects the arm $i_{t}$ with the best UCB estimate, and receives a reward $r_{t}$ corresponding to it. This is then used to update the corresponding design matrix (i.e. $\mathbf{V}_{i_{t}}$) and the estimated reward parameter vector (i.e. $\phi_{i_{t}}$) in Steps 8,9. A key difference from the previous algorithm is that in DisLinUCB the UCB bonus for arm $i\in[K]$ is computed using the arm specific design matrix i.e. $\mathbf{V}_{i}$, whereas in LinUCB   it is computed using the overall design matrix $\mathbf{M}$. In our DisLinUCB instantiation, we set the regularizer $\lambda=1$ and choose the exploration coefficient $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$. This choice of $\lambda$ and $\gamma$ is standard for LinUCB in the $d_{1}+d_{2}$ dimensional shared setting. For these values of $\lambda$ and $\gamma$, the regret analysis of LinUCB (under Assumption 2) can be appropriately modified to provide a regret guarantee of $\tilde{O}((d_{1}+d_{2})\sqrt{KT})$ for DisLinUCB. In Corollary 4, we present a much stronger guarantee when Assumption 1 is also satisfied.
Algorithm 2 DisLinUCB (Algorithm $1$ in (Li et al., 2010))
0:  Regularizer $\lambda$, exploration coefficient $\gamma$, failure probability $\delta\in(0,1)$
1:  For every arm $i\in[K]$, initialize $\phi_{i}={\mathbf{0}}\in\mathbb{R}^{d_{1}+d_{2}}$, $\mathbf{V}_{i}=\lambda\mathbf{I}_{d_{1}+d_{2}}$, $u_{i}={\mathbf{0}}\in\mathbb{R}^{d_{1}+d_{2}}$
2:  for $t=1,\dots,T$ do
3:     Receive context $\mathcal{X}_{t}=\{(x_{i},z_{i})\}_{i\in[K]}$
4:     Set $\overline{x}_{i}=\begin{bmatrix}x_{i}^{\intercal}&z_{i}^{\intercal}\end{bmatrix}^{\intercal}$ for all $i\in[K]$
5:     $i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\overline{x}_{i}},{\phi_{i}}\rangle+\gamma\left\lVert\overline{x}_{i}\right\rVert_{\mathbf{V}_{i}^{-1}}$
6:     Play arm $i_{t}$ and observe reward $r_{t}$
7:     $u_{i_{t}}\leftarrow u_{i_{t}}+r_{t}\overline{x}_{i_{t}}$
8:     $\mathbf{V}_{i_{t}}\leftarrow\mathbf{V}_{i_{t}}+\overline{x}_{i_{t}}\overline{x}_{i_{t}}^{\intercal}$
9:     Update $\phi_{i_{t}}=\mathbf{V}_{i_{t}}^{-1}u_{i_{t}}$
10:  end for
Corollary 4 (Regret of DisLinUCB).
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})KT\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.
Proof of Corollary 4 uses techniques similar to the ones used in the proof of Theorem 3 with modifications to adapt it to the disjoint setting. Complete proof is provided in Appendix C.
3.2 HyLinUCB
Our new algorithm HyLinUCB, is derived from Algorithm 1 by using different values for parameters $\lambda$ and $\gamma$. In our HyLinUCB instantiation, we set the regularizer $\lambda=K$ and choose the exploration coefficient $\gamma=2(S\sqrt{K}+\sqrt{2(d_{1}+d_{2})\log(T/\delta))}$. The rest of the algorithm is exactly the same as Algorithm 1. However, due to this change the known regret analysis of LinUCB (under Assumption 2) cannot be directly applied to get optimal regret with respect to $T$. The optimality of the guarantee relies heavily on choosing $\gamma$ correctly (since it represents confidence in estimation of model parameters) and therefore, even simple modifications do not seem to work. Our main intuition behind using this new value for $\gamma$ is that even though the feature vector post reduction to the shared setting i.e. $\tilde{x}_{i}$ (See Step $4$ in Algorithm 1) is $d_{1}+d_{2}K$ dimensional, it has only $d_{1}+d_{2}$ non-zero entries. This leads us to the question whether we can use a tighter exploration parameter that depends on the true intrinsic dimensionality of our feature vectors. In Theorem 5, we prove a strong regret guarantee (optimal with respect to $T$) when Assumptions 1 and 2 are satisfied. We also provide strong empirical evidence in support of our choice for $\gamma$ in Section 4.
Theorem 5 (Regret of HyLinUCB).
At the end of $T$ rounds, the regret of HyLinUCB (Algorithm 1 with $\lambda=K$, $\gamma=2(S\sqrt{K}+\sqrt{2(d_{1}+d_{2})\log(T/\delta)})$) under Assumptions 1 and 2 is upper bounded by
$C_{1}\sqrt{K^{3}T\log({K(d_{1}+d_{2})}/{\delta})}+C_{2}\sqrt{(d_{1}+d_{2}K)KT\log({KT(d_{1}+d_{2})}/{\delta})}$
with probability at least $1-4\delta$. Here $C_{1},C_{2}>0$ are universal constants and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
Our proof of Theorem 5 requires many new ideas on top of the techniques used in the proof of Theorem 3 in order to accommodate the new confidence parameter. Complete proof is provided in Appendix B. We finish this section with a few remarks on the algorithm.
Remark 6.
The scheme of HyLinUCB is similar to Algorithm $2$ in (Li et al., 2010). However, in (Li et al., 2010), the exploration coefficient $\gamma$ is treated as a hyperparameter to be tuned. As a result it’s extremely difficult to compute a regret guarantee and no such guarantee is known. Moreover, it also brings a practical overhead of experimentally finding the optimal parameter. However, in HyLinUCB, we fix an appropriate value of $\gamma$ and prove optimal (with respect to $T$) regret guarantees. This is where our algorithm differs from (Li et al., 2010).
Remark 7.
Line 9 in Algorithm 1 requires an inverse of a square matrix of $d_{1}+d_{2}K$ dimension, which requires $O((d_{1}+d_{2}K)^{2})$ computation if one uses the Sherman-Morrison formula (Section $0.7.4$ in Horn and Johnson (2012)). This is still quadratic in $K$. However, it can be made linear in $K$ using block matrix inversion technique (see (Li et al., 2010; Stanimirović et al., 2019)). For simplicity, we present the algorithm in the above format.
4 Experimental Setup
In this section, we provide details of our experimental setup. We perform extensive experiments under different problem settings and on real world datasets to compare HyLinUCB, LinUCB, DisLinUCB and a baseline HyRan (Kim et al., 2023). We set HyRan as our baseline because for the shared setting with similar assumptions, it was recently shown to have much better regret than all popular LinearCB baselines described in Section 1.3. Our experiments are divided into two types (a) synthetic (b) real-world. We provide further details about the setup below. The code for all the experiments are available at https://github.com/nirjhar-das/HyPay_Bandits.
4.1 Synthetic
For the synthetic experiments, the contextual arm features are generated by sampling randomly from some distribution. Since the hybrid reward setting is characterized by $3$ key parameters: $d_{1}$, $d_{2}$ and $K$, to be exhaustive in our evaluation, we design $3$ parameter settings to study the effect of these parameters on the regret of the various algorithms.
Parameter Settings:
In Setting 1, we fix $d_{1}=40$ and $d_{2}=5$ to observe the regret when $d_{2}$ is small compared to $d_{1}$. In Setting 2, we reverse the situation and fix $d_{1}=5$ and $d_{2}=40$. For both Setting 1 and Setting 2 we fix $K=25$ and $T=80,000$. For Setting 3, we fix $d_{1}=d_{2}=5$ and $T=30,000$ but vary $K$ over a set of points between $10$ and $400$. This helps us capture the dependence of regret on the number of arms $K$ for the various algorithms.
Environments:
We model the environment as a sequence of $T$ contexts (arm features) and a set of reward parameters. For each of our parameter settings, we create $5$ different environments. For each environment, we run $5$ parallel trials. In two different trials of the same environment, the sequence of contexts and the reward parameters remain unchanged, but the random noise in reward generation is allowed to vary.
Stochastic Feature Generation:
For each environment, the contexts i.e., the shared and arm-specific features of an arm in a particular round are generated by uniformly sampling $d_{1}$ and $d_{2}$ sized vectors from unit balls in $d_{1}$ and $d_{2}$ dimensions respectively.
Reward Simulation:
Reward parameters for an environment are created by uniformly sampling the shared parameter $\theta^{*}$ from the $d_{1}$-dimensional ball of radius $1$ (i.e., $S=1$). Thereafter, $K$ vectors are sampled uniformly from the $d_{2}$-dimensional ball of radius $1$, which become the disjoint parameters $\{\beta_{i}^{*}\}_{i\in[K]}$. The stochastic reward generated for arm $i\in[K]$ with features $(x_{i},z_{i})$ is
$r_{t}=\langle{x_{i}},{\theta^{*}}\rangle+\langle{z_{i}},{\beta^{*}_{i}}\rangle+\eta~{},$
where noise $\eta$ is sampled from the Gaussian distribution $\mathcal{N}(0,0.01)$.
4.2 Real-World
We also perform experiments on the Yahoo! Front Page Dataset (Yahoo! Webscope, 2010). This dataset contains click logs from real time user interactions with the news articles displayed in the Featured tab of the Today module on Yahoo! Front Page during the first $10$ days of May $2009$. This dataset was also used in (Li et al., 2010) for experiments with the hybrid model.
Dataset Description:
At every round, a user arrives and is presented with $K=20$ news articles. The user is described by $6$ features and every article presented also has $6$ features. The first feature is always $1$ while the remaining $5$ features correspond to the $5$ membership features constructed via conjoint analysis with a bi-linear model as described in (Chu et al., 2009). The dataset contains about $45$ million such user-article interaction entries. Further, one of the articles out of these $K$ articles is actually shown to the user, for which click (or no-click) is recorded.
Feature Construction:
At a particular round, suppose the user features are denoted by $u\in\mathbb{R}^{6}$ and the arm (news article) features of arm $i\in[K]$ are denoted by $v_{i}\in\mathbb{R}^{6}$. Then, for arm $i$, the shared features $x_{i}$ are given by vectorizing $uv_{i}^{\intercal}$, while the disjoint features $z_{i}$ are set equal to $v_{i}$. Thus, $d_{1}=36$ and $d_{2}=6$.
Parameter Learning:
We first learn a reward model by training a hybrid linear regression model on the first $1M$ data points (contained in May 1, 2009 dataset). For every user and the article that was actually shown to the user, we create the shared and the arm parameters $x_{i}$ and $z_{i}$ (respectively) of the arm (article). Note that the index of the arm shown in round $n$, i.e., $i_{n}$, is also an essential part of this dataset (as we need to train the arm parameters too). The target variable $y_{n}$ is the click feedback, which is a boolean variable. Thus, for $N=1M$ data points, we have $(x_{i_{n},n},z_{i_{n},n},y_{n})_{n=1}^{N}$. Finally we obtain the parameters as:
$\theta^{*},\{\beta^{*}_{i}\}_{i\in[K]}=\mathop{\mathrm{argmin}}_{\theta,\{\beta_{i}\}_{i\in[K]}}\sum_{n=1}^{N}\left(\langle{x_{i_{n},n}},{\theta}\rangle+\langle{z_{i_{n},n}},{\beta_{i_{n}}}\rangle-y_{n}\right)^{2}$
Reward Simulation:
For simulating the bandit experiments, the stochastic reward for an arm $i$ with features $x_{i}$, $z_{i}$ is generated as
$r_{t}=\langle{x_{i}},{\theta^{*}}\rangle+\langle{z_{i}},{\beta^{*}_{i}}\rangle+\eta~{},$
where noise $\eta$ is sampled from the Gaussian distribution $\mathcal{N}(0,0.0001)$. The reason for keeping the variance of the noise low is that it was observed that the mean rewards are themselves of the order $0.01$, so high noise variance causes the algorithms to take longer to demonstrate sub-linear regret due to smaller signal-to-noise ratio in each arm pull’s feedback.
5 Results
Figure 1: Results of our experiments. Top-left: Regret vs # of Rounds ($T$) for Setting 1; Top-right: Regret vs # of Rounds ($T$) for Setting 2; Bottom-left: Regret versus # of Arms for Setting 3; Bottom-right: Relative regret with respect to HyLinUCB for Yahoo! Dataset.
In all experiments, we compare HyLinUCB, LinUCB, DisLinUCB and HyRan (Kim et al., 2023).
5.1 Synthetic Experiments
Regret vs $T$:
In figure 1, in the top row, we present results for Setting 1 (left) and Setting 2 (right). We plot the cumulative regret averaged over 5 parallel trials for each of the $5$ different environments. In Setting 1, with $d_{1}\gg d_{2}$, we observe that the performance of HyLinUCB is the best followed by LinUCB, thus validating the superiority of these algorithms with higher shared parameters, which resembles more like a fully shared setting. DisLinUCB also performs comparably but the regret is higher than LinUCB. HyRan performs the worst in Setting 1 although some sub-linear nature can be observed. In Setting 2, with $d_{2}\gg d_{1}$, DisLinUCB emerges as the best algorithm, while HyLinUCB is a very close second. HyRan exhibits a linear regret in this regime and eventually crosses beyond both DisLinUCB and HyLinUCB. Notably, LinUCB has the worst performance in this setting, although the regret looks sub-linear. This matches with the intuition that Setting 2 is closer to the fully disjoint setting, hence LinUCB is not well-suited. However, the advantage of HyLinUCB is very clear from these experiments as HyLinUCB performs very well in both the settings, thus validating its suitability for the hybrid reward problem.
Regret vs $K$:
In Fig. 1, the left plot in the bottom row shows the effect of $K$ on the total regret. This plot corresponds to Setting 3, with $d_{1}=d_{2}=5$. We again perform 5 parallel trials for (each of the) 5 different environments and then calculate the average of the total regret over $T$ rounds in all these $5\times 5=25$ simulations. The X-axis in the plot represents the number of arms $K$ while the Y-axis denotes the (average) total regret. From the plot, we can observe that HyLinUCB has the smallest regret over all values of $K$ while DisLinUCB is the second best, followed by LinUCB and then HyRan. HyLinUCB displays a very slow growth with $K$, leading us to believe that its regret guarantee in Section 3 is not tight and can possibly be improved significantly.
5.2 Real-World Experiment
The results of the semi-synthetic experiment for $10M$ rounds (starting from May 2, 2009 dataset and moving to subsequent days’ dataset till $10M$ rounds are complete) is shown in Fig. 1 in the bottom right. We plot the regret of other algorithms relative to HyLinUCB (subtract the cumulative regret of HyLinUCB from the cumulative regret of the algorithm), which has the smallest regret. LinUCB comes close to HyLinUCB whereas regret of DisLinUCB and HyRan are much larger. In this experiment, we do not have any control over the context features hence it is possible that the diversity assumption is not satisfied. However, we observe that HyLinUCB still performs much better than the other algorithms, demonstrating strong evidence of good performance in hybrid models.
6 Conclusion and Future Work
In this work, we revisited the problem of linear contextual bandits with hybrid rewards which is a useful setting in many applications such as news recommendation. The problem was originally proposed in (Li et al., 2010) albeit without any theoretical guarantees. Reducing this problem to the shared and disjoint settings, we derived regret guarantee for the LinUCB and DisLinUCB (Algorithm $1$ in Li et al. (2010)) algorithms, improving over their well known guarantees under a popular diversity assumption. Finally, we propose a new algorithm HyLinUCB that employs a tighter exploration coefficient leveraging the sparsity of the problem. We also derive regret guarantee for this algorithm that has optimal dependence on $T$. We perform extensive empirical evaluation of the three algorithms in various synthetic scenarios (choices of $d_{1},d_{2}$ and $K$) as well as on real-world datasets (Yahoo! Front Page Dataset Yahoo! Webscope (2010)). We empirically compare these algorithms with the state-of-the-art HyRan (Kim et al., 2023) algorithm and demonstrate that HyLinUCB outperforms the other algorithms in almost all cases and comes very close to the best in the remaining.
Since HyLinUCB performs much better than LinUCB empirically, an interesting future direction will be to derive tighter regret bounds that will provably demonstrate the efficacy of HyLinUCB. Another future direction is to derive the regret guarantees under a different diversity assumption and understanding the trade-offs therein.
Appendix for “”
Appendix A Regret Analysis of LinUCB
Notations
We first outline the notations of this section. For a rectangular matrix $\mathbf{D}\in\mathbb{R}^{p_{1}\times p_{2}}$, we will denote by $\sigma_{max}(\mathbf{D})$ and $\sigma_{min}(\mathbf{D})$ the maximum and minimum singular values of $\mathbf{D}$. Note that singular values of $\mathbf{D}$ are given as the square root of the corresponding eigenvalue of $\mathbf{D}^{\intercal}\mathbf{D}$ (or $\mathbf{D}\mathbf{D}^{\intercal}$). Recall that in the HyLinCB problem, the learner-environment interaction over $t$ rounds gives rise to the sequence of pulled arm feature tuples $\{(x_{s},z_{s})\}_{s=1}^{t}$ and the rewards $\{r_{s}\}_{s=1}^{t}$. Note that the problem can be converted to a LinCB instance by thinking of the hybrid arm features actually coming from $\mathbb{R}^{d_{1}+d_{2}K}$ as the modified vectors $\tilde{x}_{s}=\mathcal{P}(i_{s},x_{s},z_{s})$ where $i_{s}$ is the index of the arm pulled in round $s$. We will denote by ${\mathcal{T}}_{i,t}$ as the set of rounds till $t$ that the arm $i$ gets pulled and $\tau_{i,t}=\lvert{\mathcal{T}}_{i,t}\rvert$, for all $i\in[K]$. Moreover, we will denote the failure probability with $\delta\in(0,1)$ which is a fixed constant throughout. Let us first define three key matrices we will heavily use in this section.
$\mathbf{V}_{t}\coloneqq\lambda\mathbf{I}+\sum_{s=1}^{t}x_{s}x_{s}^{\intercal}$ $\mathbf{B}_{i,t}\coloneqq\sum_{s\in{\mathcal{T}}_{i,t}}x_{s}z_{s}^{\intercal}$ $\mathbf{W}_{i,t}=\lambda\mathbf{I}+\sum_{s\in{\mathcal{T}}_{i,t}}z_{s}z_{s}^{\intercal}$
Now recall the estimators for the hybrid model $\widehat{\theta}_{t}$ and $\{\widehat{\beta}_{i,t}\}_{i=1}^{K}$ are given by:
$\begin{bmatrix}\widehat{\theta}_{t}\\ \widehat{\beta}_{1,t}\\ \vdots\\ \widehat{\beta}_{K,t}\end{bmatrix}=\mathbf{M}_{t}^{-1}\sum_{s=1}^{t}r_{s}\tilde{x}_{s}\text{ where }\mathbf{M}_{t}\coloneqq\lambda\mathbf{I}+\sum_{s=1}^{t}\tilde{x}_{s}\tilde{x}_{s}^{\intercal}=\begin{bmatrix}\mathbf{V}_{t}&\mathbf{B}_{i,t}&\dots&\mathbf{B}_{K,t}\\ \mathbf{B}^{\intercal}_{1,t}&\mathbf{W}_{1,t}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{B}^{\intercal}_{K,t}&0&\dots&\mathbf{W}_{K,t}\end{bmatrix}$ (2)
We will denote the vector of parameter estimates $\begin{bmatrix}{\widehat{\theta}_{t}}^{\intercal}&{\widehat{\beta}_{1,t}}^{\intercal}&\dots&{\widehat{\beta}_{K,t}}^{\intercal}\end{bmatrix}^{\intercal}$ with the notation $\widehat{\phi}_{t}$. Finally, we will define $\mathbf{U}_{t}$ which is the block diagonal part of $\mathbf{M}_{t}$ defined as follows:
$\mathbf{U}_{t}={\rm diag}(\mathbf{V}_{t},\mathbf{W}_{1,t},\dots\mathbf{W}_{K,t})$
A.1 Proof of Theorem 3
After round $t$, from (2) we have that $\mathbf{M}_{t}\widehat{\phi}_{t}=\sum_{s=1}^{t}r_{s}\tilde{x}_{s}$. Let ${\phi^{*}}=\begin{bmatrix}{\theta^{*}}^{\intercal}&{\beta^{*}_{1}}^{\intercal}&\dots&{\beta^{*}_{K}}^{\intercal}\end{bmatrix}^{\intercal}$. By the definition of the reward model, we have for all $s\in[T]$, $r_{s}=\langle{x_{s}},{\theta^{*}}\rangle+\langle{z_{s}},{\beta^{*}_{i_{s}}}\rangle+\eta_{s}$, where $\eta_{s}$ is the zero mean $1$-subgaussian noise and $i_{s}$ is the index of the arm pulled at round $s$.
Lemma 8.
With probability at least $1-\delta$, for all $t\geq 0$, $\left\lVert\widehat{\phi}_{t}-{\phi^{*}}\right\rVert_{\mathbf{M}_{t}}\leq\gamma\coloneqq 2\sqrt{\lambda K}S+\sqrt{2(d_{1}+d_{2}K)\log\left(T/\delta\right)}$.
Proof.
The proof is a direct application of Lemma 41 with the fact that $\left\lVert{\phi^{*}}\right\rVert_{2}=\sqrt{\left\lVert\theta^{*}\right\rVert_{2}^{2}+\sum_{i=1}^{K}\left\lVert\beta^{*}_{i}\right\rVert_{2}^{2}}\leq\sqrt{(K+1)S^{2}}=2S\sqrt{K}$. ∎
Lemma 9.
For any $\tilde{x}\in\mathbb{R}^{d_{1}+d_{2}K}$ and $t\geq 0$, with probability at least $1-\delta$, we have $\lvert\langle{\tilde{x}},{\widehat{\phi}_{t}-{\phi^{*}}}\rangle\rvert\leq\gamma\left\lVert\tilde{x}\right\rVert_{\mathbf{M}_{t}^{-1}}$.
Proof.
$\lvert\langle{\tilde{x}},{\widehat{\phi}_{t}-{\phi^{*}}}\rangle\rvert\leq\left\lVert\tilde{x}\right\rVert_{\mathbf{M}_{t}^{-1}}\cdot\left\lVert\widehat{\phi}_{t}-{\phi^{*}}\right\rVert_{\mathbf{M}_{t}}$ by Cauchy-Schwarz inequality. Thereby, applying Lemma 8 finishes the proof. ∎
We define $T_{o}$ and $T_{m}$ as: $T_{m}\coloneqq\left(\frac{16}{\rho^{2}}+\frac{8}{3\rho}\right)\log\left(\frac{2(d_{1}+d_{2})KT}{\delta}\right)$ and
$T_{o}\coloneqq\left(\frac{128}{\rho^{2}}\vee 4T_{m}\right)K^{2}\log\left(\frac{(d_{1}+d_{2})K}{\delta}\right)~{}.$
We now define the following events which will be used to prove Theorem 3.
$\mathcal{E}_{1}$$\coloneqq\left\{\forall\ i\in[K],\forall\ t\geq 0:\left\lVert\mathbf{B}_{i,t}\right\rVert\leq\sqrt{8\tau_{i,t}\log\left(K(d_{1}+d_{2})/{\delta}\right)}\right\}$
$\mathcal{E}_{2}$$\coloneqq\left\{\forall\ t\geq T_{m}:\lambda_{min}\left(\mathbf{V}_{t}\right)\geq{\rho t}/{2}\right\}$
$\mathcal{E}_{3}$$\coloneqq\left\{\forall\ i\in[K],\forall\ t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{m}:\lambda_{min}\left(\mathbf{W}_{i,t}\right)\geq{\rho\tau_{i,t}}/{2}\right\}$
$\mathcal{E}_{4}$$\coloneqq\left\{\forall\ t\geq 0:\left\lVert\widehat{\phi}_{t}-{\phi^{*}}\right\rVert_{\mathbf{M}_{t}}\leq\gamma\right\}$
Finally, we define $\mathcal{E}\coloneqq\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}\cap\mathcal{E}_{4}$
Lemma 10.
We have the following:
  1. 1.
    $\mathbb{P}[\mathcal{E}]\geq 1-4\delta$
  2. 2.
    Under event $\mathcal{E}$, for all $t\geq T_{o}$, $\frac{1}{2}\mathbf{U}_{t}\preccurlyeq\mathbf{M}_{t}\preccurlyeq 2\mathbf{U}_{t}$
Proof.
[¯$\mathbb{P}[\overline{\mathcal{E}}]=\mathbb{P}\left[(\overline{\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}})\cup\overline{\mathcal{E}_{4}}\right]\leq\mathbb{P}\left[\overline{\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}}\right]+\mathbb{P}\left[\overline{\mathcal{E}_{4}}\right]\leq 3\delta+\delta=4\delta$ where the first inequality is via Union bound and the second inequality is by application of Corollary 20 and Lemma 8. Thus, $\mathbb{P}[\mathcal{E}]\geq 1-4\delta$, which proves 1. Lastly, 2. is a restatement of Lemma 22 whose proof has been presented in Appendix A.2. ∎
Lemma 11.
Let $i_{t}$ be the arm played in round $t>T_{o}$. Then, the instantaneous regret bound of LinUCB is upper bounded by $\sqrt{8}\gamma\left(\left\lVert x_{t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ with probability at least $1-4\delta~{}.$
Proof.
Let $i^{*}_{t}$ denote the best arm in round $t$. Hereon, we will assume that the event $\mathcal{E}$ holds, which happens with probability at least $1-4\delta~{}.$ For clarity, let $\widehat{\phi}=\widehat{\phi}_{t-1}$ The instantaneous regret in round $t$ is given by
$\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$ $\leq\langle{\tilde{x}_{i^{*}_{t},t}},{\widehat{\phi}}\rangle+\gamma\left\lVert\tilde{x}_{i^{*}_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}-\langle{x_{i_{t},t}},{\widehat{\phi}}\rangle+\gamma\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}$ (by Lemma 9)
$\leq\langle{\tilde{x}_{i_{t},t}},{\widehat{\phi}}\rangle+\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}-\langle{x_{i_{t},t}},{\widehat{\phi}}\rangle+\gamma\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}$ ($i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\tilde{x}_{i,t}},{\widehat{\phi}}\rangle+\gamma\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}$)
$=2\gamma\left\lVert\tilde{x}_{i_{t}}\right\rVert_{\mathbf{M}_{t-1}^{-1}}$
$\leq 2\sqrt{2}\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{U}_{t-1}^{-1}}$ (by Lemma 10)
$=\sqrt{8}\gamma\sqrt{\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}^{2}+\left\lVert z_{i_{t}}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}^{2}}$
$\leq\sqrt{8}\gamma\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ ($\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for $a,b\geq 0$)
Lemma 12.
Under event $\mathcal{E}$, we have the following:
$\sum_{t=T_{o}+1}^{T}\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ $\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}~{}.$
Proof.
The proof of this Lemma is deferred till Appendix A.2
Theorem 13 (Regret of LinUCB).
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=2S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)KT\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
Proof.
The total regret after $T$ rounds is the sum of the instantaneous regrets over these rounds. Suppose $T>T_{o}$ (it is easy to see that a loose lower bound on $T_{o}$ is $\tilde{\Omega}(K^{4})$). Hereon, we will assume that the event $\mathcal{E}$ holds, which happens with probability at least $1-4\delta~{}.$ Hence,
$\texttt{Reg}(T,\texttt{LinUCB})$
$=\sum_{t=1}^{T}\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$
$\leq 2\gamma\sum_{t=1}^{T_{o}}\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{t-1}}+\sqrt{8}\gamma\sum_{t=T_{o}+1}^{T}\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ (by Lemma 11)
$\leq 2\gamma\sqrt{4(d_{1}+d_{2}K)T_{o}\log(T_{o})}+\sqrt{8}\gamma\sum_{t=T_{o}+1}^{T}\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ (Lemma 45 on first term)
$\leq 4\gamma\sqrt{(d_{1}+d_{2}K)T_{o}\log(T_{o})}+\sqrt{8}\gamma\left(2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}\right)$ (Lemma 12)
Finally, since we have $\gamma=2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$, substituting in the above expression,
$\texttt{Reg}(T,\texttt{LinUCB})$ $\leq c_{1}\left(2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}\right)$
$+16\left(\frac{\sqrt{K}}{\sqrt{\rho}}\right)\left(2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}\right)\sqrt{T}$
where $c_{1}=4\sqrt{(d_{1}+d_{2}K)T_{o}\log(T_{o})}+4K\sqrt{2d_{2}T_{m}\log(T_{m})}$. Simplifying constants and only keeping the asymptotically dominating terms in $T$,
$\texttt{Reg}(T,\texttt{LinUCB})\leq\frac{C}{\sqrt{\rho}}\sqrt{(d_{1}+d_{2}K)KT\log(T/\delta)}$
where $C>0$ is a universal constant. ∎
A.2 Supporting Lemmas for A.1
Proposition 14.
$\mathbf{M}_{t}\left(\widehat{\phi}_{t}-{\phi^{*}}\right)=\sum_{s=1}^{t}\eta_{s}\tilde{x}_{s}-\lambda{\phi^{*}}$ for all $t\geq 0$.
Proof.
The mean reward of an arm $i\in[K]$ can also be written as $\langle{\tilde{x}_{i}},{{\phi^{*}}}\rangle$. Note that $\mathbf{M}_{t}=\lambda\mathbf{I}+\sum_{s=1}^{t}\tilde{x}_{s}\tilde{x}_{s}^{\intercal}$. Consequently,
$\mathbf{M}_{t}{\phi^{*}}=\lambda{\phi^{*}}+\sum_{S=1}^{t}\langle{\tilde{x}_{s}},{{\phi^{*}}}\rangle\tilde{x}_{s}$
Plugging this in the expression $\mathbf{M}_{t}\left(\widehat{\phi}_{t}-{\phi^{*}}\right)$ and using (2), we get:
$\mathbf{M}_{t}\left(\widehat{\phi}_{t}-{\phi^{*}}\right)$ $=-\lambda{\phi^{*}}+\sum_{s=1}^{t}r_{s}\tilde{x}_{s}-\sum_{s=1}^{t}\langle{\tilde{x}_{s}},{{\phi^{*}}}\rangle\tilde{x}_{s}$
$=-\lambda{\phi^{*}}+\sum_{s=1}^{t}\left(r_{s}-\langle{\tilde{x}_{s}},{{\phi^{*}}}\rangle\right)\tilde{x}_{s}=\sum_{s=1}^{t}\eta_{s}\tilde{x}_{s}-\lambda{\phi^{*}}$
Proposition 15.
$\mathbf{U}^{-\frac{1}{2}}_{t}\mathbf{M}_{t}\mathbf{U}_{t}^{-\frac{1}{2}}$ $=\mathbf{I}+\begin{bmatrix}{\mathbf{0}}&\mathbf{V}_{t}^{-\frac{1}{2}}\mathbf{B}_{1,t}\mathbf{W}_{1,t}^{-\frac{1}{2}}&\dots&\mathbf{V}_{t}^{-\frac{1}{2}}\mathbf{B}_{K,t}\mathbf{W}_{K,t}^{-\frac{1}{2}}\\ \mathbf{W}_{1,t}^{-\frac{1}{2}}\mathbf{B}_{1,t}^{\intercal}\mathbf{V}_{t}^{-\frac{1}{2}}&{\mathbf{0}}&\dots&{\mathbf{0}}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{W}_{K,t}^{-\frac{1}{2}}\mathbf{B}_{K,t}^{\intercal}\mathbf{V}_{t}^{-\frac{1}{2}}&{\mathbf{0}}&\dots&{\mathbf{0}}\end{bmatrix}$
$\equiv\mathbf{I}+\mathbf{A}_{t}$ (definition of $\mathbf{A}_{t}$)
In what follows, we show that the maximum eigenvalue of $\mathbf{A}_{t}$ is bounded away from $1$ under the assumption 1, which in turn would imply Lemma 22.
Lemma 16.
Let $\mathbf{Z}_{t}=\begin{bmatrix}\mathbf{V}_{t}^{-\frac{1}{2}}\mathbf{B}_{1,t}\mathbf{W}_{1,t}^{-\frac{1}{2}}&\dots&\mathbf{V}_{t}^{-\frac{1}{2}}\mathbf{B}_{K,t}\mathbf{W}_{K,t}^{-\frac{1}{2}}\end{bmatrix}$. Then, it holds that almost surely, $\lambda_{max}(\mathbf{A}_{t})=-\lambda_{min}(\mathbf{A}_{t})=\sigma_{max}(\mathbf{Z}_{t})$.
Proof.
First, note that $\mathbf{A}_{t}=\begin{bmatrix}{\mathbf{0}}&\mathbf{Z}_{t}\\ \mathbf{Z}_{t}^{\intercal}&{\mathbf{0}}\end{bmatrix}$. Thus, $\mathbf{A}_{t}$ is the Hermitian dilation (see def. 43) of $\mathbf{Z}_{t}$. Thus, by Lemma 44, the square of eigenvalues of $\mathbf{A}_{t}$ are equal to the square of the singular values of $\mathbf{Z}_{t}$. Consequently, for every positive singular value $\sigma$ of $\mathbf{Z}_{t}$, there are two eigenvalues of $\mathbf{A}_{t}$ namely, $\sigma$ and $-\sigma$. Hence, clearly, $\lambda_{max}(\mathbf{A}_{t})=\sigma_{max}(\mathbf{Z}_{t})$ and $\lambda_{min}(\mathbf{A}_{t})=-\sigma_{max}(\mathbf{Z}_{t})$. ∎
Now we turn to upper bounding $\left\lVert\mathbf{Z}_{t}\right\rVert$. Our goal is to show that $\left\lVert\mathbf{Z}_{t}\right\rVert$ is bounded away from $1$. First, we describe two high probability events for $t$ sufficiently large.
Lemma 17.
Let $\delta\in(0,1)$. Fix an $i\in[K]$. Then, for all $t\geq 0$, $\left\lVert\mathbf{B}_{i,t}\right\rVert\leq\sqrt{8\tau_{i,t}\log((d_{1}+d_{2})K/\delta)}$ with probability at least $1-\frac{\delta}{K}$.
Proof.
We have $\mathbf{B}_{i,t}=\sum_{s\in{\mathcal{T}}_{i}}x_{s}z_{s}^{\intercal}$. By assumption 1, we have $\mathbb{E}_{s-1}\left[x_{s}z_{s}^{\intercal}\right]={\mathbf{0}}~{}.$ Hence, $\mathbf{B}_{i,t}$ is a sum of Martingale difference sequences. Moreover, since $\left\lVert x_{s}\right\rVert_{2}\leq 1$ and $\left\lVert z_{s}\right\rVert_{2}\leq 1$, for the Hermitian dilation (definition 43) $\mathbf{H}_{s}$ of $x_{s}z_{s}^{\intercal}$,
$\mathbf{H}_{s}^{2}=\begin{bmatrix}\left\lVert z_{s}\right\rVert_{2}^{2}x_{s}x_{s}^{\intercal}&{\mathbf{0}}\\ {\mathbf{0}}&\left\lVert x_{s}\right\rVert_{2}^{2}z_{s}z_{s}^{\intercal}\end{bmatrix}\preccurlyeq\mathbf{I}_{d_{1}+d_{2}}$
Thus, by Matrix Azuma inequality (Lemma 39), we can upper bound the operator norm of $\mathbf{B}_{i,t}$ as follows:
$\mathbb{P}\left[\exists t\geq 0:\left\lVert\mathbf{B}_{i,t}\right\rVert\geq\varepsilon\right]\leq(d_{1}+d_{2})\exp\left(-\frac{\varepsilon^{2}}{8\tau_{i,t}^{2}}\right)$
Setting $\varepsilon=\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}$, we obtain,
$\mathbb{P}\left[\exists t\geq 0:\left\lVert\mathbf{B}_{i,t}\right\rVert\geq\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}\right]\leq\frac{\delta}{K}$
Corollary 18.
For all $i\in[K]$ and $t\geq 0$, $\left\lVert\mathbf{B}_{i,t}\right\rVert\leq\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}$ with probability at least $1-\delta~{}.$
Proof.
Proof follows by applying union bound on Lemma 17. ∎
Hereon, let us define the event
$\mathcal{E}_{1}\coloneqq\left\{\forall\ i\in[K],t\geq 0,\ \left\lVert\mathbf{B}_{i,t}\right\rVert\leq\sqrt{8\tau_{i,t}\log\left(\frac{K(d_{1}+d_{2})}{\delta}\right)}\right\}$
Thus, we have that $\mathbb{P}[\mathcal{E}_{1}]\geq 1-\delta$.
Next, we describe another event that happens with high probability under assumption 1.
Lemma 19.
We first define the quantity $T_{m}\coloneqq\left(\frac{16}{\rho^{2}}+\frac{8}{3\rho}\right)\log\left(\frac{2(d_{1}+d_{2})KT}{\delta}\right)$. Then, $\mathbb{P}\left[\forall t\geq T_{m}:\lambda_{min}(\mathbf{V}_{t})\geq\frac{\rho}{2}t\right]\geq 1-\delta$, and $\mathbb{P}\big{[}\forall i\in[K],\tau_{i,t}\geq T_{m}:\lambda_{min}(\mathbf{W}_{i,t})\geq\frac{\rho}{2}\tau_{i,t}\big{]}\geq 1-\delta$.
Proof.
The result is obtained by directly applying Lemma 38 followed by union bound. ∎
Hereon, we will denote by $\mathcal{E}_{2}$ as the event: $\{\forall\ i\in[K],\ \forall\ \tau_{i,t}\geq T_{m},\lambda_{min}(\mathbf{W}_{i,t})\geq\rho\tau_{i,t}/2\}$ and by $\mathcal{E}_{3}$ the event: $\{\forall\ t\geq T_{m},\lambda_{min}(\mathbf{V}_{t})\geq\rho t/2\}$. Finally, we will denote
$\mathcal{E}_{o}\coloneqq\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$
Corollary 20.
$\mathbb{P}[\mathcal{E}_{o}]\geq 1-3\delta~{}.$
Proof.
By Union Bound, $\mathbb{P}[\overline{\mathcal{E}_{o}}]\leq\mathbb{P}[\overline{\mathcal{E}_{1}}]+\mathbb{P}[\overline{\mathcal{E}_{2}}]+\mathbb{P}[\overline{\mathcal{E}_{3}}]\leq 3\delta~{}.$
Lemma 21.
Under event $\mathcal{E}_{o}$, for all $t>T_{o}\coloneqq\left(\frac{128}{\rho^{2}}\vee 4T_{m}\right)K^{2}\log\left(\frac{(d_{1}+d_{2})K}{\delta}\right)$, we have $\left\lVert\mathbf{Z}_{t}\right\rVert\leq\frac{1}{2}$
Proof.
Firstly, note that by definition, $\left\lVert\mathbf{Z}_{t}\right\rVert=\sup_{\left\lVert b\right\rVert_{2}=1}\left\lVert\mathbf{Z}_{t}b\right\rVert_{2}$. Hence,
$\left\lVert\mathbf{Z}_{t}\right\rVert$ $=\sup_{\left\lVert b\right\rVert_{2}\leq 1}\left\lVert\mathbf{Z}_{t}b\right\rVert_{2}$
$=\sup_{\sum_{i=1}^{K}\left\lVert b_{i}\right\rVert_{2}^{2}\leq 1}\left\lVert\sum_{i=1}^{K}\mathbf{V}_{t}^{-1/2}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-1/2}b_{i}\right\rVert_{2}$
$\leq\sup_{\sum_{i=1}^{K}\left\lVert b_{i}\right\rVert_{2}^{2}\leq 1}\sum_{i=1}^{K}\left\lVert\mathbf{V}_{t}^{-1/2}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-1/2}b_{i}\right\rVert_{2}$
$\leq\sum_{i=1}^{K}\sup_{\left\lVert b_{i}\right\rVert_{2}\leq 1}\left\lVert\mathbf{V}_{t}^{-1/2}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-1/2}b_{i}\right\rVert_{2}$
$=\sum_{i=1}^{K}\left\lVert\mathbf{V}_{t}^{-1/2}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-1/2}\right\rVert$
$\leq\sum_{i=1}^{K}\left\lVert\mathbf{V}_{t}^{-1/2}\right\rVert\left\lVert\mathbf{B}_{i,t}\right\rVert\left\lVert\mathbf{W}_{i,t}^{-1/2}\right\rVert$
$=\sum_{i=1}^{K}\frac{\left\lVert\mathbf{B}_{i,t}\right\rVert}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}$ ($\lambda_{\max}\left(\mathbf{D}^{-1}\right)=\frac{1}{\lambda_{min}(\mathbf{D})}$)
$\leq\sum_{i=1}^{K}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}$ (3)
We can split the RHS in (3) into two parts, one in which the arms have been pulled for less than $T_{m}$ times and the another containing the rest of the arms. Thus, we have
$\left\lVert\mathbf{Z}_{t}\right\rVert\leq\sum_{i:\tau_{i,t}<T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}+\sum_{i:\tau_{i,t}\geq T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}$ (4)
We first bound the second term in (4) as follows
$\sum_{i:\tau_{i,t}\geq T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}$ $\leq\sum_{i:\tau_{i,t}\geq T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{(\rho t/2)}\sqrt{(\rho\tau_{i,t}/2)}}$ ($\mathcal{E}$ holds and $T_{m}\leq\tau_{i,t}$)
$=\sum_{i:\tau_{i,t}\geq T_{m}}\frac{4\sqrt{2\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\rho^{2}t}}$
$\leq\sum_{i:\tau_{i,t}\geq T_{m}}\frac{1}{2K}$ ($t\geq\frac{128}{\rho^{2}}K^{2}\log\left(\frac{(d_{1}+d_{2})K}{\delta}\right)$)
Next, the first term in (4) can be bounded as:
$\sum_{i:\tau_{i,t}<T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\lambda_{min}(\mathbf{V}_{t})}\sqrt{\lambda_{min}(\mathbf{W}_{i,t})}}$ $\leq\sum_{i:\tau_{i,t}<T_{m}}\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\rho t/2}\sqrt{\lambda}}$ ($\mathcal{E}$ holds; $\lambda_{min}(\mathbf{W}_{i,t})\geq\lambda$)
$\leq\sum_{i:\tau_{i,t}<T_{m}}\frac{4\sqrt{T_{m}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\rho\lambda t}}$
$\leq\sum_{i:\tau_{i,t}<T_{m}}\frac{1}{\sqrt{\lambda\rho}}\cdot\frac{1}{2K}$ ($t\geq 4K^{2}T_{m}\log(K(d_{1}+d_{2})/\delta)$)
Choosing $\lambda\geq\frac{1}{\rho}$, we can have the RHS above smaller than $\frac{1}{2K}$. Combining these results, we finally have,
$\left\lVert\mathbf{Z}_{t}\right\rVert\leq\sum_{i:\tau_{i,t}<T_{m}}\frac{1}{2K}+\sum_{i:\tau_{i,t}\geq T_{m}}\frac{1}{2K}=\frac{1}{2}$
Lemma 22.
Under event $\mathcal{E}_{o}$, for all $t\geq T_{o}$, $\frac{1}{2}\mathbf{U}_{t}\preccurlyeq\mathbf{M}_{t}\preccurlyeq 2\mathbf{U}_{t}$.
Proof.
We first note that by Lemma 16, $\lambda_{max}(\mathbf{A}_{t})=-\lambda_{min}(\mathbf{A}_{t})$. Thus, under $\mathcal{E}_{o}$, $\lambda_{max}(\mathbf{A}_{t})=\left\lVert\mathbf{Z}_{t}\right\rVert\leq\frac{1}{2}$ implies that $\lambda_{min}(\mathbf{A}_{t})\geq-\frac{1}{2}$. Thus, all eigenvalues of $\mathbf{A}_{t}$ lie in the range $[-1/2,1/2]$. This in turn implies that the eigenvalues of $\mathbf{I}+\mathbf{A}_{t}$ lie in the range $[1/2,3/2]$. From this, it can be concluded that,
$\frac{1}{2}\mathbf{I}\preccurlyeq\mathbf{I}+\mathbf{A}_{t}\preccurlyeq 2\mathbf{I}$
as eigenvalues of $2\mathbf{I}$ dominate that of $\mathbf{I}+\mathbf{A}_{t}$ which in turn dominate that of $(1/2)\mathbf{I}$. Thus, we have, by combining with Proposition 15, $\frac{1}{2}\mathbf{I}\preccurlyeq\mathbf{U}_{t}^{-\frac{1}{2}}\mathbf{M}_{t}\mathbf{U}_{t}^{-\frac{1}{2}}\preccurlyeq 2\mathbf{I}$, which then implies that,
$\frac{1}{2}\mathbf{U}_{t}\preccurlyeq\mathbf{M}_{t}\preccurlyeq 2\mathbf{U}_{t}$
See 12
Proof.
For a given $i\in[K]$, let the elements in ${\mathcal{T}}_{i,T}$ be rearranged so that they are in increasing order. Let ${\mathcal{T}}_{i,\xi}$ denote the subset of ${\mathcal{T}}_{i,T}$ that contains only the first $T_{m}$ elements if $\left|{\mathcal{T}}_{i,T}\right|>T_{m}$ else ${\mathcal{T}}_{i,\xi}={\mathcal{T}}_{i,T}$. Finally, let $\tau_{i,\xi}=\left|{\mathcal{T}}_{i,\xi}\right|$. Hence, the second term above can now be written as
$\sum_{t=T_{o}}^{T}\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}$ $\leq\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}}\left\lVert z_{i,s}\right\rVert_{\mathbf{W}_{i,s-1}^{-1}}$
$=\sum_{i=1}^{K}\left(\sum_{s\in{\mathcal{T}}_{i,\xi}}\left\lVert z_{i,s}\right\rVert_{\mathbf{W}_{i,s-1}^{-1}}+\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert z_{i,s}\right\rVert_{\mathbf{W}_{i,s-1}^{-1}}\right)$
$\leq\sum_{i=1}^{K}\sqrt{4d_{2}\tau_{i,\xi}\log(\tau_{i,\xi})}+\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert z_{i,s}\right\rVert_{\mathbf{W}_{i,s-1}^{-1}}$ (by Lemma 45)
$\leq\sum_{i=1}^{K}\sqrt{4d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert z_{i,s}\right\rVert_{\mathbf{W}_{i,s-1}^{-1}}$ ($\tau_{i,\xi}\leq T_{m}$)
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\frac{\left\lVert z_{i,s}\right\rVert_{2}}{\sqrt{\lambda_{min}(\mathbf{W}_{i,s-1})}}$
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\frac{1}{\sqrt{\rho(\tau_{i,s}-1)/2}}$ (under $\mathcal{E}$, $\lambda_{min}(\mathbf{W}_{i,t})\geq\frac{\rho\tau_{i,t}}{2}$ since $\tau_{i,t}\geq T_{m}$; $\left\lVert z_{i,s}\right\rVert_{2}\leq 1$)
$=2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\mathbb{I}\left\{\tau_{i,T}>T_{m}\right\}\sum_{s=T_{m}}^{\tau_{i,T}}\frac{1}{\sqrt{\rho(s-1)/2}}$
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\mathbb{I}\left\{\tau_{i,T}>T_{m}\right\}\sqrt{\frac{2}{\rho}}\int_{s=1}^{\tau_{i,T}+1}\frac{1}{\sqrt{s}}$
$=2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\mathbb{I}\left\{\tau_{i,T}>T_{m}\right\}\sqrt{\frac{8\tau_{i,T}}{\rho}}$
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sum_{i=1}^{K}\sqrt{\frac{8\tau_{i,T}}{\rho}}$
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{8}{\rho}}\sqrt{K\cdot\sum_{i=1}^{K}\tau_{i,T}}$ (Cauchy-Schwarz)
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{8KT}{\rho}}$ ($\sum_{i=1}^{K}\tau_{i,T}=T$)
Fnally, we have,
$\sum_{t=T_{o}}^{T}\left\lVert x_{i_{t}}\right\rVert_{\mathbf{V}_{t-1}^{-1}}$ $\leq\sum_{t=T_{o}}^{T}\frac{\left\lVert x_{i_{t}}\right\rVert_{2}}{\sqrt{\lambda_{min}(\mathbf{V}_{t-1})}}\leq\sum_{t=T_{o}}^{T}\frac{1}{\sqrt{\rho(t-1)/2}}$ (by Lemma 10, $\lambda_{min}(\mathbf{V}_{t})\geq\rho t/2$; $\left\lVert x_{i_{t},t}\right\rVert_{2}\leq 1$)
$\leq\sqrt{\frac{2}{\rho}}\int_{t=1}^{T+1}\frac{1}{\sqrt{t}}\leq\sqrt{\frac{8T}{\rho}}$
Putting things back,
$\sum_{t=T_{o}}^{T}\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)$ $\leq\sqrt{\frac{8T}{\rho}}+2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{8KT}{\rho}}$
$\leq 2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}$
Appendix B Regret Analysis of HyLinUCB
In this section, we will use the same notations as specified in Appendix A.1 except, in this section
$\gamma=2(\sqrt{2\lambda}S+\sqrt{2(d_{1}+d_{2})\log\frac{T}{\delta}})~{}.$
B.1 Proof of Theorem 5
Lemma 23.
Under event $\mathcal{E}$ (defined in Appendix A.1), for $t>T_{o}$, for all $i\in[K]$, $\lvert\langle{\tilde{x}_{i,t}},{\widehat{\phi}_{t-1}-{\phi^{*}}}\rangle\rvert\leq 2(\sqrt{2\lambda}S+\sqrt{2(d_{1}+d_{2})\log\frac{T}{\delta}})\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+O(\frac{1}{\sqrt{t-1}})~{}.$
Proof.
Let us denote by $\Delta\widehat{\phi}_{t-1}=\widehat{\phi}_{t-1}-{\phi^{*}}$. We can write,
$\mathbf{U}_{t-1}\Delta\widehat{\phi}_{t-1}$ $=\mathbf{M}_{t-1}\Delta\widehat{\phi}_{t-1}+(\mathbf{U}_{t-1}-\mathbf{M}_{t-1})\Delta\widehat{\phi}_{t-1}$
$=\mathbf{M}_{t-1}\Delta\widehat{\phi}_{t-1}-\mathbf{U}_{t-1}^{1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}$ (by Proposition 15)
Thus, $\Delta\widehat{\phi}_{t-1}=\mathbf{U}_{t-1}^{-1}\mathbf{M}_{t-1}\Delta\widehat{\phi}_{t-1}-\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}$. Consequently,
$\langle{\tilde{x}_{i,t}},{\Delta\widehat{\phi}_{t-1}}\rangle$ $=\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1}\mathbf{M}_{t-1}\Delta\widehat{\phi}_{t-1}}\rangle-\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle$
$=\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1}\left(\sum_{s=1}^{t-1}\eta_{s}\tilde{x}_{s}-\lambda{\phi^{*}}\right)}\rangle-\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle$
$=\langle{\mathbf{U}_{t-1}^{-1}\tilde{x}_{i,t}},{\sum_{s=1}^{t-1}\eta_{s}\tilde{x}_{s}-\lambda{\phi^{*}}}\rangle-\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle$ (5)
Since $\mathbf{U}_{t-1}$ is block diagonal, $\mathbf{U}^{-1}_{t-1}={\rm diag}(\mathbf{V}_{t-1}^{-1},\mathbf{W}_{1,t-1}^{-1},\dots,\mathbf{W}_{K,t-1}^{-1})$. Also, $\tilde{x}_{i,t}$ is a sparse vector, with zeros everywhere except first $d_{1}$ coordinates and coordinates starting from $d_{1}+(i-1)d_{2}+1$ till $d_{1}+id_{2}$. Thus, $\mathbf{U}_{t-1}^{-1}\tilde{x}_{i,t}=\begin{bmatrix}(\mathbf{V}_{t-1}^{-1}x_{i,t})^{\intercal}&{\mathbf{0}}&\dots&(\mathbf{W}_{i,t-1}^{-1}z_{i,t})^{\intercal}&\dots&{\mathbf{0}}\end{bmatrix}^{\intercal}$. Therefore,
$\langle{\mathbf{U}_{t-1}^{-1}\tilde{x}_{i,t}},{\sum_{s=1}^{t-1}\eta_{s}\tilde{x}_{s}-\lambda{\phi^{*}}}\rangle$ $=\langle{\mathbf{V}_{t-1}^{-1}x_{i,t}},{\sum_{s=1}^{t-1}\eta_{s}x_{s}-\lambda\theta^{*}}\rangle$
$\quad\quad+\langle{\mathbf{W}_{i,t-1}^{-1}z_{i,t}},{\sum_{s\in{\mathcal{T}}_{i,t-1}}\eta_{s}z_{s}-\lambda\beta^{*}_{i}}\rangle~{}.$
Plugging this in eq. (5),
$\left|\langle{\tilde{x}_{i,t}},{\Delta\widehat{\phi}_{t-1}}\rangle\right|$
$=\left\lvert\langle{\mathbf{V}_{t-1}^{-1}x_{i,t}},{\sum_{s=1}^{t-1}\eta_{s}x_{s}-\lambda\theta^{*}}\rangle+\langle{\mathbf{W}_{i,t-1}^{-1}z_{i,t}},{\sum_{s\in{\mathcal{T}}_{i,t-1}}z_{s}-\lambda\beta^{*}_{i}}\rangle\right.$
$\quad\quad\quad\left.-\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\right\rvert$
$\leq\lvert\langle{\mathbf{V}_{t-1}^{-1}x_{i,t}},{\sum_{s=1}^{t-1}\eta_{s}x_{s}-\lambda\theta^{*}}\rangle\rvert+\lvert\langle{\mathbf{W}_{i,t-1}^{-1}z_{i,t}},{\sum_{s\in{\mathcal{T}}_{i,t-1}}z_{s}-\lambda\beta^{*}_{i}}\rangle\rvert$
$\quad\quad\quad+\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$
$\leq\left\lVert x_{i,t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}\cdot\left\lVert\sum_{s=1}^{t-1}\eta_{s}x_{s}-\lambda\theta^{*}\right\rVert_{\mathbf{V}_{t-1}}+\left\lVert z_{i,t}\right\rVert_{\mathbf{W}_{i,t-1}^{-1}}\cdot\left\lVert\sum_{s\in{\mathcal{T}}_{i,t-1}}\eta_{s}z_{s}-\lambda\beta^{*}_{i}\right\rVert_{\mathbf{W}_{i,t-1}}$
$\quad\quad\quad+\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$
$\leq(\sqrt{2d_{1}\log(T/\delta)}+\sqrt{\lambda}S)\left\lVert x_{i,t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+(\sqrt{2d_{2}\log(\tau_{i,t-1}/\delta)}+\sqrt{\lambda}S)\left\lVert z_{i,t}\right\rVert_{\mathbf{W}_{i,t-1}^{-1}}$
$\quad\quad\quad+\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$ (by Lemma 40)
Next, we upper bound $\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$ as follows:
$\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$ $\leq\lambda_{max}(\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2})\left\lVert\tilde{x}_{i,t}\right\rVert_{2}\left\lVert\Delta\widehat{\phi}_{t-1}\right\rVert_{2}$ (Sub-multiplicativity of norm)
$\leq\lambda_{max}(\mathbf{A}_{t-1})\cdot S\cdot\left\lVert\Delta\widehat{\phi}_{t-1}\right\rVert_{2}$ (Lemma 46; Assumption 2)
From proof of Lemma 21, it can be extracted that,
$\lambda_{max}(\mathbf{A}_{t})\leq 4K\sqrt{\log(K(d_{1}+d_{2})/\delta)}\left(\frac{\sqrt{T_{m}}}{\sqrt{\rho\lambda t}}+\frac{\sqrt{2}}{\sqrt{\rho^{2}t}}\right)$
Next, we have that $\left\lVert\Delta\widehat{\phi}_{t-1}\right\rVert_{2}\leq\frac{\left\lVert\Delta\widehat{\phi}_{t-1}\right\rVert_{\mathbf{M}_{t-1}}}{\sqrt{\lambda}}$ since $\lambda_{min}(\mathbf{M}_{t-1})\geq\lambda$. Lastly, via Lemma 8, $\left\lVert\Delta\widehat{\phi}_{t-1}\right\rVert_{\mathbf{M}_{t-1}}\leq 2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(\frac{T}{\delta})}$. Putting everything back together,
$\lvert\langle{\tilde{x}_{i,t}},{\mathbf{U}_{t-1}^{-1/2}\mathbf{A}_{t-1}\mathbf{U}_{t-1}^{1/2}\Delta\widehat{\phi}_{t-1}}\rangle\rvert$
$\leq 4KS\sqrt{\log(\frac{K(d_{1}+d_{2})}{\delta})}\left(\frac{\sqrt{T_{m}}}{\sqrt{\rho\lambda t}}+\frac{\sqrt{2}}{\sqrt{\rho^{2}t}}\right)\frac{2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(\frac{T}{\delta})}}{\sqrt{\lambda}}$
$=\frac{\alpha(\lambda)}{\sqrt{t}}$ (definition of $\alpha(\lambda)$)
Thus, we have,
$\left|\langle{\tilde{x}_{i,t}},{\Delta\widehat{\phi}_{t-1}}\rangle\right|\leq\left(\sqrt{2d_{1}\log(\frac{T}{\delta})}+\sqrt{\lambda}S\right)\left\lVert x_{i,t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\left(\sqrt{2d_{2}\log(\frac{\tau_{i,t-1}}{\delta})}+\sqrt{\lambda}S\right)\left\lVert z_{i,t}\right\rVert_{\mathbf{W}_{i,t-1}^{-1}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$
$\leq\sqrt{2\left(2(d_{1}+d_{2})\log(\frac{T}{\delta})+2\lambda S^{2}\right)\cdot\left(\left\lVert x_{i,t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}^{2}+\left\lVert z_{i,t}\right\rVert_{\mathbf{W}_{i,t-1}^{-1}}^{2}\right)}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$ (Cauchy-Schwarz)
$=\sqrt{2\left(2(d_{1}+d_{2})\log(\frac{T}{\delta})+2\lambda S^{2}\right)\cdot\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{U}_{t-1}^{-1}}^{2}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$
$\leq\sqrt{2\left(2(d_{1}+d_{2})\log(\frac{T}{\delta})+2\lambda S^{2}\right)\cdot 2\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}^{2}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$ (Lemma 22)
$\leq 2\left(\sqrt{2(d_{1}+d_{2})\log(\frac{T}{\delta})}+S\sqrt{2\lambda}\right)\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$
Lemma 24.
Under event $\mathcal{E}$, after $T_{o}$ rounds have elapsed, the instantaneous regret of HyLinUCB is upper bounded as:
$\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle\leq 2\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\frac{2\alpha(\lambda)}{\sqrt{t-1}}$
Proof.
We have the following:
$\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle$ $\leq\langle{\tilde{x}_{i^{*}_{t},t}},{\widehat{\phi}_{t-1}}\rangle+\gamma\left\lVert\tilde{x}_{i^{*}_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$
$\leq\langle{\tilde{x}_{i_{t},t}},{\widehat{\phi}_{t-1}}\rangle+\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\frac{\alpha(\lambda)}{\sqrt{t-1}}$ (6)
where the second inequality is because $i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\tilde{x}_{i,t}},{\widehat{\phi}_{t-1}}\rangle+\gamma\left\lVert\tilde{x}_{i,t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}$. Similarly, we have:
$\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$ $\geq\langle{\tilde{x}_{i_{t},t}},{\widehat{\phi}_{t-1}}\rangle-\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}-\frac{\alpha(\lambda)}{\sqrt{t-1}}$ (7)
Using eq. 6 and 7, $\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle\leq 2\gamma\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\frac{2\alpha(\lambda)}{\sqrt{t-1}}$
Now we restate Theorem 5 and present its proof.
Theorem 25 (Regret of HyLinUCB).
At the end of $T$ rounds, the regret of HyLinUCB (Algorithm 1 with $\lambda=K$, $\gamma=\sqrt{KS}+\sqrt{2(d_{1}+d_{2})\log(T/\delta)}$) under Assumptions 1, 2 is upper bounded by
$C_{1}\sqrt{K^{3}T\log({K(d_{1}+d_{2})}/{\delta})}+C_{2}\sqrt{(d_{1}+d_{2}K)KT\log({KT(d_{1}+d_{2})}/{\delta})}$
with probability at least $1-4\delta$. Here $C_{1},C_{2}>0$ are universal constants and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
Proof.
The proof is similar to the proof of Theorem 13. We will assume throughout that event $\mathcal{E}$ holds, which happens with probability at least $1-4\delta$ (Lemma 10). Let $i^{*}_{t}$ and $i_{t}$ be the best arm and the arm played in round $t$ respectively. Then, following similar arguments as in the proof of Theorem 13, we have,
$\texttt{Reg}(T,\texttt{HyLinUCB})=\sum_{t=1}^{T}\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$
$=\sum_{t=1}^{T_{o}}\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle+\sum_{t=T_{o}+1}^{T}\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$
$\leq 2T_{o}S+\sum_{t=T_{o}}^{T}\langle{\tilde{x}_{i^{*}_{t},t}},{{\phi^{*}}}\rangle-\langle{\tilde{x}_{i_{t},t}},{{\phi^{*}}}\rangle$ (rewards are bounded in $[-S,S]$)
$\leq 2T_{o}S+2\gamma\sum_{t=T_{o}+1}^{T}\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{M}_{t-1}^{-1}}+\sum_{t=T_{o}+1}^{T}\frac{2\alpha(\lambda)}{\sqrt{t-1}}$ (Lemma 24)
$\leq 2T_{o}S+2\sqrt{2}\gamma\sum_{t=T_{o}+1}^{T}\left\lVert\tilde{x}_{i_{t},t}\right\rVert_{\mathbf{U}_{t-1}^{-1}}+\sum_{t=T_{o}+1}^{T}\frac{2\alpha(\lambda)}{\sqrt{t-1}}$ (Lemma 22)
$=2T_{o}S+\sqrt{8}\gamma\sum_{t=T_{o}+1}^{T}\left(\left\lVert x_{i_{t},t}\right\rVert_{\mathbf{V}_{t-1}^{-1}}+\left\lVert z_{i_{t},t}\right\rVert_{\mathbf{W}_{i_{t},t-1}^{-1}}\right)+\sum_{t=T_{o}+1}^{T}\frac{2\alpha(\lambda)}{\sqrt{t-1}}$
$\leq 2T_{o}S+\sqrt{8}\gamma\left(2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}\right)+\sum_{t=T_{o}+1}^{T}\frac{2\alpha(\lambda)}{\sqrt{t-1}}$ (Lemma 12)
$\leq 2T_{o}S+\sqrt{8}\gamma\left(2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}\right)+\int_{t=1}^{T+1}\frac{2\alpha(\lambda)}{\sqrt{t-1}}$
$\leq 2T_{o}S+\sqrt{8}\gamma\left(2K\sqrt{d_{2}T_{m}\log(T_{m})}+\sqrt{\frac{32KT}{\rho}}\right)+4\alpha(\lambda)\sqrt{T}$
Finally, from definition of $\alpha(\lambda)$, one can observe that setting $\lambda=K$, and letting $c=4S\sqrt{\log(K(d_{1}+d_{2})/\delta)}$ and $\beta=2S\sqrt{\lambda K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$
$\alpha(\lambda)$ $=c\beta K\left(\frac{\sqrt{T_{m}}}{\lambda\sqrt{\rho}}+\frac{\sqrt{2}}{\rho}\frac{1}{\sqrt{\lambda}}\right)$
$=c\left(\beta\sqrt{\frac{T_{m}}{\rho}}+\frac{\sqrt{2}}{\rho}\left(2SK^{3/2}+\sqrt{2(d_{1}+d_{2}K)K\log(T/\delta)}\right)\right)$
$\leq\frac{C_{o}}{\rho}\left(S^{2}K^{3/2}+S\sqrt{(d_{1}+d_{2}K)K\log(T/\delta)}\right)\sqrt{\log(K(d_{1}+d_{2})/\delta)}$
for some universal constant $C_{o}$. Therefore, keeping only asymptotically dominating terms in $T$ and $K$,
$\texttt{Reg}(T,\texttt{HyLinUCB})$ $\leq C_{1}\sqrt{K^{3}T\log(\frac{K(d_{1}+d_{2})}{\delta})}$
$\quad\quad+C_{2}\sqrt{(d_{1}+d_{2}K)KT\log(\frac{KT(d_{1}+d_{2})}{\delta})}$
where $C_{1},C_{2}>0$ are universal constants. ∎
Appendix C Regret Analysis of DisLinUCB
In the algorithm of DisLinUCB, since the parameters for each arm are estimated in a disjoint fashion, let us denote (with a slight abuse of notation) by $\phi^{*}_{i}\coloneqq\begin{bmatrix}{\theta^{*}}^{\intercal}&{\beta^{*}_{i}}^{\intercal}\end{bmatrix}^{\intercal}$ the true parameters that determine the reward of every arm $i\in[K]$. Moreover, we overload the notation $\widehat{\phi}$ to denote by $\widehat{\phi}_{i}\coloneqq\begin{bmatrix}{\widehat{\theta}}^{\intercal}_{i}&{\widehat{\beta}_{i}}^{\intercal}\end{bmatrix}^{\intercal}$ the estimate of $\phi^{*}_{i}$ for every arm $i\in[K]$. For every round $t\in[T]$ and every arm $i\in[K]$, let $\overline{x}_{i,t}\coloneqq\begin{bmatrix}x_{i,t}^{\intercal}&z_{i,t}^{\intercal}\end{bmatrix}^{\intercal}$ be the feature vector of arm $i$ in round $t$ formed by appending the shared and the action features. Finally, we define $\mathbf{M}_{i,t}\coloneqq\sum_{s\in{\mathcal{T}}_{i,t}}\overline{x}_{i,s}\overline{x}_{i,s}^{\intercal}+\lambda\mathbf{I}$ for this section. Since DisLinUCB also solves a ridge regression to obtain the parameter estimates, we have, after round $t$,
$\widehat{\phi}_{i,t}=\mathbf{M}_{i,t}^{-1}\sum_{s\in{\mathcal{T}}_{i,t}}r_{s}\overline{x}_{s},\quad\forall\ i\in[K]$ (8)
C.1 Proof of Corollary 4 from Section 3.1
First, we will present the key lemmas that will be used to prove Corollary 4. Then the detailed proof is presented which follows the same recipe used in proving Theorem 13.
Proposition 26.
For all $i\in[K]$ and $t\geq 0$, $\mathbf{M}_{i,t}\left(\widehat{\phi}_{i,t}-\phi^{*}_{i}\right)=\sum_{s\in{\mathcal{T}}_{i,t}}\eta_{s}\overline{x}_{s}-\lambda\phi^{*}_{i}$
Lemma 27.
With probability at least $1-\delta$, for all $i\in[K]$ and $t\geq 0$, we have that $\left\lVert\widehat{\phi}_{i,t}-\phi^{*}_{i}\right\rVert_{\mathbf{M}_{i,t}}\leq\gamma\coloneqq 2\sqrt{\lambda S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$
Proof.
The proof is a direct application of Lemma 41 with the fact that $\left\lVert\phi^{*}_{i}\right\rVert_{2}=\sqrt{\left\lVert\theta^{*}\right\rVert_{2}^{2}+\left\lVert\beta^{*}_{i}\right\rVert_{2}^{2}}\leq\sqrt{2S^{2}}=\sqrt{2}S\leq 2\sqrt{S}$. ∎
Lemma 28.
For any given $i\in[K]$ and $t\geq 0$, with probability at least $1-\delta$, we have $\left|\langle{\overline{x}_{i,t}},{\widehat{\phi}_{i,t}-\phi^{*}_{i}}\rangle\right|\leq\gamma\left\lVert\overline{x}_{i,t}\right\rVert_{\mathbf{M}_{i,t}^{-1}}$
Proof.
$\lvert\langle{\tilde{x}},{\widehat{\phi}_{t}-{\phi^{*}}}\rangle\rvert\leq\left\lVert\tilde{x}\right\rVert_{\mathbf{M}_{t}^{-1}}\cdot\left\lVert\widehat{\phi}_{t}-{\phi^{*}}\right\rVert_{\mathbf{M}_{t}}$ by Cauchy-Schwarz inequality. Thereby, applying Lemma 8 finishes the proof. ∎
Recall that $T_{m}=\left(\frac{16}{\rho^{2}}+\frac{8}{3\rho}\right)\log\left(\frac{2(d_{1}+d_{2})KT}{\delta}\right)$ and define
$T^{\prime}_{o}\coloneqq\big{(}\frac{128}{\rho^{2}}\log(K(d_{1}+d_{2})/\delta)\big{)}\vee T_{m}~{}.$
Next, we define the following events which will be used in the regret analysis of DisLinUCB:
$\mathcal{E}_{1}$ =$=\left\{\forall\ i\in[K],\forall\ t\geq 0,\text{ s.t. }\tau_{i,t}\geq T^{\prime}_{o}:\lambda_{min}\left(\mathbf{M}_{i,t}\right)\geq\rho\tau_{i,t}/4\right\}$
$\mathcal{E}_{2}$ =$=\left\{\forall\ i\in[K],\forall\ t\geq 0,\left\lVert\widehat{\phi}_{i,t}-\phi^{*}_{i}\right\rVert_{\mathbf{M}_{i,t}}\leq\gamma\right\}$
$\mathcal{E}$ $=\mathcal{E}_{1}\cap\mathcal{E}_{2}$
Lemma 29.
$\mathbb{P}[\mathcal{E}]\geq 1-4\delta$
Proof.
By Corollary 37, $\mathbb{P}[\mathcal{E}_{1}]\geq 1-3\delta$ and by Lemma 27, $\mathbb{P}[\mathcal{E}_{2}]\geq 1-\delta$. Thus, by Union Bound, $\mathbb{P}[\mathcal{E}]\geq 1-4\delta$. ∎
Now, we restate Corollary 4 and present its proof.
Corollary 30 (Regret of DisLinUCB).
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})KT\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.
Proof.
Throughout the proof, we will assume that event $\mathcal{E}$ holds, which happens with probability at least $1-4\delta$. Let $i^{*}_{t}$ denote the best arm in round $t$. Further, to simplify notation, let $\widehat{\phi}_{i,t-1}$ be written as $\widehat{\phi}_{i}$ for all $i\in[K]$. The instantaneous regret in round $t$ is upper bounded as:
$\langle{\overline{x}_{i^{*}_{t},t}},{\phi^{*}_{i^{*}_{t}}}\rangle-\langle{\overline{x}_{i_{t},t}},{\phi^{*}_{i_{t}}}\rangle$
$\leq\langle{\overline{x}_{i^{*}_{t},t}},{\widehat{\phi}_{i^{*}_{t}}}\rangle+\gamma\left\lVert\overline{x}_{i^{*}_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i^{*}_{t},t-1}}-\langle{\overline{x}_{i_{t},t}},{\widehat{\phi}_{i_{t}}}\rangle+\gamma\left\lVert\overline{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i_{t},t-1}}$ (Lemma 28)
$\leq\langle{\overline{x}_{i_{t},t}},{\widehat{\phi}_{i_{t}}}\rangle+\gamma\left\lVert\overline{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i_{t},t-1}}-\langle{\overline{x}_{i_{t},t}},{\widehat{\phi}_{i_{t}}}\rangle+\gamma\left\lVert\overline{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i_{t},t-1}}$ ($i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\overline{x}_{i,t}},{\widehat{\phi}_{i}}\rangle+\gamma\left\lVert\overline{x}_{i,t}\right\rVert_{\mathbf{M}^{-1}_{i,t-1}}$)
$=2\gamma\left\lVert\overline{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i_{t},t-1}}$
Thus, the cumulative regret of DisLinUCB over $T$ rounds can be upper bounded by:
$\texttt{Reg}(T,\texttt{DisLinUCB})\leq\sum_{t=1}^{T}2\gamma\left\lVert\overline{x}_{i_{t},t}\right\rVert_{\mathbf{M}^{-1}_{i_{t},t-1}}=2\gamma\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$ (9)
Now, for a given $i\in[K]$, let the elements in ${\mathcal{T}}_{i,T}$ be rearranged so that they are in increasing order. Let ${\mathcal{T}}_{i,\xi}$ denote the subset of ${\mathcal{T}}_{i,T}$ that contains only the first $T^{\prime}_{o}$ elements if $\left|{\mathcal{T}}_{i,T}\right|>T^{\prime}_{o}$ else ${\mathcal{T}}_{i,\xi}={\mathcal{T}}_{i,T}$. Finally, let $\tau_{i,\xi}=\left|{\mathcal{T}}_{i,\xi}\right|$. Thus, we can write the sum in the RHS above as:
$\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,T}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$ $=\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}+\sum_{i:\tau_{i,T}>T_{o}}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$
To upper bound the first term above, we will use Elliptic Potential Lemma (Lemma 45) but for the second term above, since the matrices $\mathbf{M}_{i,s}$ contains the sum of at least $T_{o}$ rank one (of the type $\overline{x}\overline{x}^{\intercal}$) terms, we can apply the eigenvalue guarantee of event $\mathcal{E}$. Specifically,
$\sum_{i=1}^{K}\sum_{s\in{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$
$\leq\sum_{i=1}^{K}\sqrt{2(d_{1}+d_{2})\left|{\mathcal{T}}_{i,\xi}\right|\log\left(\left|{\mathcal{T}}_{i,\xi}\right|\right)}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$ (Lemma 45 on first term)
$\leq\sum_{i=1}^{K}\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\sum_{s\in{\mathcal{T}}_{i,T}\setminus{\mathcal{T}}_{i,\xi}}\left\lVert\overline{x}_{i,s}\right\rVert_{\mathbf{M}^{-1}_{i,s-1}}$ ($\left|{\mathcal{T}}_{i,\xi}\right|<T^{\prime}_{o}\ \forall\ i\in[K]$)
$\leq K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\sum_{s=T^{\prime}_{o}+1}^{\tau_{i,T}}\frac{2\sqrt{2}}{\sqrt{\rho(s-1)}}$ ($\lambda_{min}\left(\mathbf{M}_{i,s}\right)\geq\rho\tau_{i,s}/4$ under $\mathcal{E}$; $\left\lVert\overline{x}_{i,s}\right\rVert_{2}\leq\sqrt{2}$)
$\leq K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\int_{s=1}^{\tau_{i,T}+1}\frac{2\sqrt{2}}{\sqrt{\rho(s-1)}}$
$=K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\sum_{i:\tau_{i,T}>T^{\prime}_{o}}\frac{4\sqrt{2}}{\sqrt{\rho}}\sqrt{\tau_{i,T}}$
$\leq K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\frac{4\sqrt{2}}{\sqrt{\rho}}\sum_{i=1}^{K}\sqrt{\tau_{i,T}}$
$\leq K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\frac{4\sqrt{2}}{\sqrt{\rho}}\sqrt{K\sum_{i=1}^{K}\tau_{i,T}}$ (Cauchy-Schwarz)
$=K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\frac{4\sqrt{2}}{\sqrt{\rho}}\sqrt{KT}$
Thus, by plugging this upper bound back in eq. (9), we obtain,
$\texttt{Reg}(T,\texttt{DisLinUCB})\leq 2\gamma\left(K\sqrt{2(d_{1}+d_{2})T^{\prime}_{o}\log\left(T^{\prime}_{o}\right)}+\frac{4\sqrt{2}}{\sqrt{\rho}}\sqrt{KT}\right)$
Noting that $\gamma=O(\sqrt{(d_{1}+d_{2})}\log(KT/\delta))$ and writing only the asymptotically dominating term, we have, for a universal constant $C>0$,
$\texttt{Reg}(T,\texttt{DisLinUCB})\leq C\sqrt{(d_{1}+d_{2})KT\log(KT/\delta)}$
C.2 Supporting Lemmas for Appendix C.1
Let us define $\mathbf{U}_{i,t}=\begin{bmatrix}\mathbf{V}_{i,t}&\mathbf{B}_{i,t}\\ \mathbf{B}^{\intercal}_{i,t}&\mathbf{W}_{i,t}\end{bmatrix}$ where $\mathbf{V}_{i,t}=\lambda\mathbf{I}+\sum_{s\in{\mathcal{T}}_{i,t}}x_{i,s}x_{i,s}^{\intercal}$, $\mathbf{W}_{i,t}=\lambda\mathbf{I}+\sum_{s\in{\mathcal{T}}_{i,t}}z_{i,s}z_{i,s}^{\intercal}$ and $\mathbf{B}_{i,t}=\sum_{s\in{\mathcal{T}}_{i,t}}x_{i,s}z_{i,s}^{\intercal}$ for all $i\in[K]$.
Proposition 31.
For any $i\in[K]$ and $t\geq 0$,
$\mathbf{U}_{i,t}^{-\frac{1}{2}}\mathbf{M}_{i,t}\mathbf{U}_{i,t}^{-\frac{1}{2}}=\mathbf{I}+\begin{bmatrix}{\mathbf{0}}&\mathbf{V}_{i,t}^{-\frac{1}{2}}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-\frac{1}{2}}\\ \mathbf{W}_{i,t}^{-\frac{1}{2}}\mathbf{B}_{i,t}^{\intercal}\mathbf{V}_{i,t}^{-\frac{1}{2}}&{\mathbf{0}}\end{bmatrix}\equiv\mathbf{I}+\mathbf{A}_{i,t}$ (definition of $\mathbf{A}_{i,t}$)
Proof.
The proof is by directly substituting $\mathbf{U}_{i,t}$ and $\mathbf{M}_{i,t}$ and noting that since $\mathbf{U}_{i,t}$ is a block diagonal matrix, $\mathbf{U}_{i,t}^{-\frac{1}{2}}={\rm diag}\left(\mathbf{V}_{i,t}^{-\frac{1}{2}},\mathbf{W}_{i,t}^{-\frac{1}{2}}\right)$. ∎
Lemma 32.
Let $\mathbf{Z}_{i,t}=\mathbf{V}_{i,t}^{-\frac{1}{2}}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-\frac{1}{2}}$ for all $i\in[K]$ and $t\geq 0$. Then, almost surely, $\lambda_{max}(\mathbf{A}_{i,t})=-\lambda_{min}(\mathbf{A}_{i,t})=\sigma_{max}(\mathbf{Z}_{i,t})$.
Proof.
The proof follows the same set of arguments presented in the proof of Lemma 16. Specifically, since $\mathbf{A}_{i,t}=\begin{bmatrix}{\mathbf{0}}&\mathbf{Z}_{i,t}\\ \mathbf{Z}_{i,t}^{\intercal}&{\mathbf{0}}\end{bmatrix}$ is the Hermitian dilation of $\mathbf{Z}_{i,t}$, thus by Lemma 44, the square of eigenvalues of $\mathbf{A}_{i,t}$ are equal to the square of the singular values of $\mathbf{Z}_{i,t}$. Consequently, for every positive singular value $\sigma$ of $\mathbf{Z}_{i,t}$, there are two eigenvalues of $\mathbf{A}_{i,t}$ namely, $\sigma$ and $-\sigma$. Hence, clearly, $\lambda_{max}(\mathbf{A}_{i,t})=\sigma_{max}(\mathbf{Z}_{i,t})$ and $\lambda_{min}(\mathbf{A}_{i,t})=-\sigma_{max}(\mathbf{Z}_{i,t})$. ∎
Now, we will turn to upper bounding $\left\lVert\mathbf{Z}_{i,t}\right\rVert$ for all $i\in[K]$.
Hereon, let us define the event
$\mathcal{E}_{1}\coloneqq\left\{\forall\ i\in[K],t\geq 0,\ \left\lVert\mathbf{B}_{i,t}\right\rVert\leq\sqrt{8\tau_{i,t}\log\left(\frac{K(d_{1}+d_{2})}{\delta}\right)}\right\}$
Thus, from Corollary 18 in Appendix A.2, we have that $\mathbb{P}[\mathcal{E}_{1}]\geq 1-\delta$.
Lemma 33.
We first define the quantity $T_{m}\coloneqq\left(\frac{16}{\rho^{2}}+\frac{8}{3\rho}\right)\log\left(\frac{2(d_{1}+d_{2})KT}{\delta}\right)$. Then, $\mathbb{P}\left[\forall i\in[K],t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{m},\lambda_{min}(\mathbf{V}_{i,t})\geq\frac{\rho}{2}\tau_{i,t}\right]\geq 1-\delta$, and $\mathbb{P}\big{[}\forall i\in[K],t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{m},\ \lambda_{min}(\mathbf{W}_{i,t})\geq\frac{\rho}{2}\tau_{i,t}\big{]}\geq 1-\delta$.
Proof.
The result is obtained by directly applying Lemma 38 followed by union bound. ∎
$\mathcal{E}_{2}$ =$=\{\forall i\in[K],t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{m},\lambda_{min}(\mathbf{V}_{i,t})\geq\frac{\rho}{2}\tau_{i,t}\}$
$\mathcal{E}_{3}$ =$=\{\forall i\in[K],t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{m},\ \lambda_{min}(\mathbf{W}_{i,t})\geq\frac{\rho}{2}\tau_{i,t}\}$
$\mathcal{E}_{o}$ $\coloneqq\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{3}$ (10)
Corollary 34.
$\mathbb{P}[\mathcal{E}_{o}]\geq 1-3\delta~{}.$
Proof.
By Union Bound, $\mathbb{P}[\overline{\mathcal{E}_{o}}]\leq\mathbb{P}[\overline{\mathcal{E}_{1}}]+\mathbb{P}[\overline{\mathcal{E}_{2}}]+\mathbb{P}[\overline{\mathcal{E}_{3}}]\leq 3\delta~{}.$
Lemma 35.
Under event $\mathcal{E}_{o}$, if for a given $i\in[K]$, $\tau_{i,t}\geq T_{o}\coloneqq\big{(}\frac{128}{\rho^{2}}\log(K(d_{1}+d_{2})/\delta)\big{)}\vee T_{m}$, then $\sigma_{max}\left(\mathbf{Z}_{i,t}\right)\leq\frac{1}{2}$.
Proof.
$\left\lVert\mathbf{Z}_{i,t}\right\rVert$ $=\left\lVert\mathbf{V}_{i,t}^{-\frac{1}{2}}\mathbf{B}_{i,t}\mathbf{W}_{i,t}^{-\frac{1}{2}}\right\rVert$
$\leq\left\lVert\mathbf{V}_{i,t}^{-\frac{1}{2}}\right\rVert\left\lVert\mathbf{B}_{i,t}\right\rVert\left\lVert\mathbf{W}_{i,t}^{-\frac{1}{2}}\right\rVert$ (Sub-multiplicativity of norm)
$=\frac{\left\lVert\mathbf{B}_{i,t}\right\rVert}{\sqrt{\lambda_{min}\left(\mathbf{V}_{i,t}\right)}\sqrt{\lambda_{min}\left(\mathbf{W}_{i,t}\right)}}$
$\leq\frac{\sqrt{8\tau_{i,t}\log(K(d_{1}+d_{2})/\delta)}}{\sqrt{\rho\tau_{i,t}/2}\sqrt{\rho\tau_{i,t}/2}}$ (Event $\mathcal{E}_{o}$ holds, $\tau_{i,t}\geq T_{m}$)
$=\sqrt{\frac{32\log(K(d_{1}+d_{2})/\delta)}{\rho^{2}\tau_{i,t}}}$
$\leq\frac{1}{2}$ ($\tau_{i,t}\geq 128\log(K(d_{1}+d_{2})/\delta)/\rho^{2}$)
Lemma 36.
Under event $\mathcal{E}_{o}$, for all $i\in[K]$ such that $\tau_{i,t}\geq T_{o}$, $\frac{1}{2}\mathbf{U}_{i,t}\preccurlyeq\mathbf{M}_{i,t}\preccurlyeq 2\mathbf{U}_{i,t}$.
Proof.
We first note that by Lemma 16, $\lambda_{max}(\mathbf{A}_{i,t})=-\lambda_{min}(\mathbf{A}_{i,t})$. Thus, under $\mathcal{E}_{o}$ and $\tau_{i,t}\geq T_{o}$, $\lambda_{max}(\mathbf{A}_{i,t})=\left\lVert\mathbf{Z}_{i,t}\right\rVert\leq\frac{1}{2}$ implies that $\lambda_{min}(\mathbf{A}_{i,t})\geq-\frac{1}{2}$. Thus, all eigenvalues of $\mathbf{A}_{i,t}$ lie in the range $[-1/2,1/2]$. This in turn implies that the eigenvalues of $\mathbf{I}+\mathbf{A}_{i,t}$ lie in the range $[1/2,3/2]$. From this, it can be concluded that,
$\frac{1}{2}\mathbf{I}\preccurlyeq\mathbf{I}+\mathbf{A}_{i,t}\preccurlyeq 2\mathbf{I}$
as eigenvalues of $2\mathbf{I}$ dominate that of $\mathbf{I}+\mathbf{A}_{i,t}$ which in turn dominate that of $(1/2)\mathbf{I}$. Thus, we have, by combining with Proposition 31, $\frac{1}{2}\mathbf{I}\preccurlyeq\mathbf{U}_{i,t}^{-\frac{1}{2}}\mathbf{M}_{i,t}\mathbf{U}_{i,t}^{-\frac{1}{2}}\preccurlyeq 2\mathbf{I}$, which then implies that,
$\frac{1}{2}\mathbf{U}_{i,t}\preccurlyeq\mathbf{M}_{i,t}\preccurlyeq 2\mathbf{U}_{i,t}$
Corollary 37.
$\mathbb{P}\left[\forall i\in[K],t\geq 0\text{ s.t. }\tau_{i,t}\geq T_{o}:\lambda_{min}(\mathbf{M}_{i,t})\geq\rho\tau_{i,t}/4\right]\geq 1-3\delta$.
Proof.
Assume that event $\mathcal{E}_{o}$ holds, which happens with probability $1-3\delta$ (Corollary 34). Then, by Lemma 36, $\mathbf{M}_{i,t}\succcurlyeq\frac{1}{2}\mathbf{U}_{i,t}$. Thus, $\lambda_{min}(\mathbf{M}_{i,t})\geq\lambda_{min}(\frac{1}{2}\mathbf{U}_{i,t})=\frac{1}{2}\lambda_{min}({\rm diag}(\mathbf{V}_{i,t},\mathbf{W}_{i,t}))=\frac{1}{2}(\lambda_{min}(\mathbf{V}_{i,t})\wedge\lambda_{min}(\mathbf{W}_{i,t}))\geq\frac{\rho\tau_{i,t}}{4}$. ∎
Appendix D Concentration Results of Random Matrices and Vectors
Lemma 38 (Lemma 7 of [Chatterji et al., 2020]).
Suppose $\{x_{s}\}_{s=1}^{t}$ is a stochastic process in $\mathbb{R}^{d}$ such that for the filtration $\mathcal{F}_{t}=\{x_{s}\}_{s=1}^{t}$, it holds that $\mathbb{E}\left[x_{t}\mid\mathcal{F}_{t-1}\right]={\mathbf{0}}$ and $\mathbb{E}\left[x_{t}x_{t}^{\intercal}\mid\mathcal{F}_{t-1}\right]\succcurlyeq\rho\mathbf{I}$. Moreover, suppose $\left\lVert x_{s}\right\rVert_{2}\leq 1$ for all $s\geq 1$. Define the matrix $\mathbf{V}_{t}$ as
$\mathbf{V}_{t}:=\mathbf{I}+\sum_{s=1}^{t}x_{s}x_{s}^{\intercal}.$
Then, with probability at least $1-\delta$, we have
$\lambda_{min}(\mathbf{V}_{t})\geq 1+\frac{\rho t}{2}$
for all $\left(16/\rho^{2}+8/(3\rho)\right)\log\left(2dT/\delta\right)\leq t\leq T$.
Lemma 39 (Matrix Azuma (Theorem 7.1 [Tropp, 2012]).
Let $\{\mathbf{X}_{k}\}_{k=1}^{\infty}$ be a matrix martingale difference sequence in $\mathbb{R}^{d_{1}\times d_{2}}$ and let $\mathcal{H}(\mathbf{X}_{k})$ be its Hermitian Dilation (definition 43). Further, let $\{\mathbf{A}_{k}\}_{k=1}^{\infty}$ be a fixed sequence of matrices in $\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$ such that,
$\mathbb{E}_{k-1}[\mathbf{X}_{k}]={\mathbf{0}}\quad\text{and}\quad\mathcal{H}(\mathbf{X}_{k})^{2}\preccurlyeq\mathbf{A}_{k}^{2}\quad\text{almostsurely.}$.
Further, let $\sigma^{2}_{t}=\lambda_{max}\left(\sum_{k=1}^{t}\mathbf{A}_{k}^{2}\right)$ for $t\geq 1$. Then, for all $\varepsilon\geq 0$,
$\mathbb{P}\left[\exists t\geq 1:\sigma_{max}\left(\sum_{k=1}^{t}\mathbf{X}_{k}\right)\geq\varepsilon\right]\leq(d_{1}+d_{2})\exp\left(-\frac{\varepsilon^{2}}{8\sigma^{2}_{t}}\right)$
Lemma 40 (Theorem 1 of Abbasi-Yadkori et al. [2011]).
Let $\{\mathcal{F}_{t}\}_{t=1}^{\infty}$ be a filtration. Let $\{x_{t}\}_{t=1}^{\infty}$ be a stochastic process in $\mathbb{R}^{d}$ with $\left\lVert x_{t}\right\rVert_{2}\leq 1$ such that $x_{t}$ is $\mathcal{F}_{t}$ measurable. Let $\{\eta_{t}\}_{t=1}^{\infty}$ be a real-valued stochastic process such that $\eta_{t}$ is $\mathcal{F}_{t}$-measurable and $\eta_{t}$ is conditionally $1$-sub-Gaussian, i.e., $\mathbb{E}[\exp(\ell\eta_{t})\mid\mathcal{F}_{t}]\leq\exp(\ell^{2}/2)$ for all $\ell\in\mathbb{R}$. Let $\lambda>0$ and for any $t\geq 1$ define:
$S_{t}=\sum_{s=1}^{t}\eta_{s}x_{s}\quad\quad\quad\mathbf{H}_{t}=\sum_{s=1}^{t}x_{s}x_{s}^{\intercal}+\lambda\mathbf{I}$
Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all $0\leq t\leq T$
$\left\lVert S_{t}\right\rVert_{\mathbf{H}^{-1}_{t}}\leq\sqrt{d\log\left(\frac{1+t/d\lambda}{\delta}\right)}\leq\sqrt{2d\log(T/\delta)}$
Lemma 41 (Theorem 2 of Abbasi-Yadkori et al. [2011]).
Let $\widehat{\theta}$ be the least squares estimate $\mathbf{H}_{t}^{-1}\left(\sum_{s=1}^{t}y_{s}x_{s}\right)$ over the feature-target pairs $\{(x_{s},y_{s})\}_{s=1}^{t}$, where $y_{s}=\langle{x_{s}},{\theta^{*}}\rangle+\eta_{s}$, and $\eta_{s}$ and $\mathbf{H}_{t}$ is as defined in Lemma 40. Let $\left\lVert\theta^{*}\right\rVert\leq S$ and $\left\lVert x_{s}\right\rVert\leq 1$ for all $s\in[t]$. Then, for all $0\leq t\leq T$, with probability at least $1-\delta$,
$\left\lVert\widehat{\theta}-\theta^{*}\right\rVert_{\mathbf{H}_{t}}\leq\sqrt{\lambda}S+\sqrt{2d\log(T/\delta)}$
Appendix E Some Useful Results
Lemma 42 [Horn and Johnson, 2012, Section 0.8.5]).
Let $\mathbf{M}=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix}$ be a block matrix with invertible $\mathbf{A}$ and $\mathbf{D}$. Then,
$\det(\mathbf{M})=\det(\mathbf{A})\det(\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B})$
Definition 43 (Hermitian Dilation).
The Hermitian dilation of a rectangular matrix $\mathbf{B}\in\mathbb{R}^{d_{1}\times d_{2}}$ is defined as the Hermitian matrix
$\mathbf{H}=\begin{bmatrix}{\mathbf{0}}&\mathbf{B}\\ \mathbf{B}^{\intercal}&{\mathbf{0}}\end{bmatrix}$
Note that $\mathbf{H}^{2}=\begin{bmatrix}\mathbf{B}\mathbf{B}^{\intercal}&{\mathbf{0}}\\ {\mathbf{0}}&\mathbf{B}^{\intercal}\mathbf{B}\end{bmatrix}~{}.$
Lemma 44 ([Tropp, 2015, Eq. (2.1.28)]).
The squared singular values of a rectangular matrix $\mathbf{B}$ coincides with the squared eigenvalues of its Hermitian dilation $\mathbf{H}$.
Lemma 45 (Elliptic Potential Lemma  [Abbasi-Yadkori et al., 2011, Lemma 10]).
Let $x_{1},x_{2},\dots x_{t}$ be a sequence of vectors in ${\mathbb{R}}^{d}$ and let $\left\lVert x_{s}\right\rVert_{2}\leq 1$ for all $s\in[t]$. Further, let $\mathbf{V}_{s}=\sum_{m=1}^{s-1}x_{m}x_{m}^{\intercal}+\lambda\mathbf{I}$. Suppose $\lambda\geq 1$. Then,
$\sum_{s=1}^{t}\left\lVert x_{s}\right\rVert_{\mathbf{V}_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{t}{\lambda d}\right)\leq 4d\log(t)$ (11)
Lemma 46 ([Horn and Johnson, 2012, Theorem 1.3.3]).
Suppose $\mathbf{A}$ is an invertible and $\mathbf{X}$ is a symmetric matrix. Then, eigenvalues of $\mathbf{A}^{-1}\mathbf{X}\mathbf{A}$ are the same as eigenvalues of $\mathbf{X}$. Particularly, $\lambda_{max}\left(\mathbf{A}^{-1}\mathbf{X}\mathbf{A}\right)=\lambda_{max}\left(\mathbf{X}\right)$.
Appendix F Empirical Validation of Assumption 1
Although the previous works make this assumption for the completely shared setting, which is indeed algorithm dependent, there are problem instances in the shared setting that allow all algorithms to satisfy this assumption (see [Papini et al., 2021, Section 3]). However, to verify whether the diversity assumption is a valid assumption in the hybrid setting, we perform experiments with algorithms LinUCB, HyLinUCB and DisLinUCB. The key implication of these assumptions on which the improved regret guarantees of HyLinUCB and LinUCB rely is that the minimum eigenvalue of the following matrices should grow linearly: (i) $\mathbf{V}_{t}=\mathbf{I}+\sum_{s=1}^{t}x_{i_{s},s}x_{i_{s},s}^{\intercal}$ and (ii) $\mathbf{W}_{i,t}=\mathbf{I}+\sum_{s=1}^{t}z_{i_{s},s}z_{i_{s},s}^{\intercal}\mathbb{I}\left\{i_{s}=i\right\}$. Besides, the maximum singular value of $\mathbf{B}_{i,t}=\sum_{s=1}^{t}x_{i_{s},s}z_{i_{s},s}^{\intercal}\mathbb{I}\left\{i_{s}=i\right\}$ for all $i\in[K]$ should grow as $\sqrt{\tau_{i,t}}$ (the number of times arm $i$ is pulled till time $t$) for HyLinUCB and LinUCB. Lastly, for DisLinUCB, this assumption’s key implication is that the minimum eigenvalue of $\mathbf{I}+\sum_{s=1}^{t}\overline{x}_{i_{s},s}\overline{x}_{i_{s},s}^{\intercal}$ grows linearly with $\tau_{i,t}$, where $\overline{x}_{i_{s},s}=\begin{bmatrix}x_{i_{s},s}^{\intercal}&z_{i_{s},s}^{\intercal}\end{bmatrix}^{\intercal}$.
Setting $d_{1}=d_{2}=10$, $K=25$ and $T=5000$, we randomly sample a realization of the reward parameters $\theta^{*}$ and $\{\beta^{*}_{i}\}_{i\in[K]}$ uniformly from a unit ball of appropriate dimensions, and then perform $N=100$ trials of the algorithms. The arm features are also generated by uniformly sampling from the respective unit balls in every round in every trial. The results are presented in Fig. 2. From the result, it can be noted that (i) the minimum eigenvalue of the relevant matrices grows linearly with rounds (or number of arm pulls) for all the algorithms and (ii) the maximum singular value of the relevant matrices grow as $\sqrt{t}$ with round $t$ in both LinUCB and HyLinUCB. These are the chief implications of Assumption 1 used in the proofs of the regret guarantees. The experimental validation therefore provides strong grounds for the diversity assumption.
Figure 2: Empirical validation of the key implications of Assumption 1. From top to bottom: LinUCB, HyLinUCB and DisLinUCB. In LinUCB: top plot shows minimum eigenvalue of $\mathbf{V}_{t}$; bottom left plot shows minimum eigenvalue of $\mathbf{W}_{i,t}$; bottom right shows maximum singular value of $\mathbf{B}_{i,t}$. Similar scheme is followed for HyLinUCB. In DisLinUCB, only the maximum eigenvalue of the relevant matrix is shown.